### 27 Mar 2026 v1

# Detection Is Harder Than Estimation in Certain Regimes: Inference for Moment and Cumulant Tensors

Abstract
We study estimation and detection of high-order moment and cumulant tensors from $n$ i.i.d. observations of a $p$-dimensional random vector, with performance measured in tensor spectral norm. On the statistical side, we prove that under sub-Gaussianity, the minimax rate for estimating the order-$d$ moment and cumulant tensors is $\sqrt{p/n}\wedge 1$. In contrast to covariance estimation, the sample moment tensor is generally no longer rate-optimal for higher-order moments. We therefore develop an estimator that attains the minimax rate up to logarithmic factors through a convex feasibility formulation over an $\varepsilon$-net of the unit sphere. On the computational side, we study the problem of testing whether the $d$-th order cumulant tensor vanishes after whitening. Using the low-degree polynomial framework, we provide evidence that detection is computationally hard when $n\ll p^{d/2}$. At the same time, we identify a regime in which an efficiently computable estimator attains error smaller than the separation at which low-degree tests can reliably distinguish the null from the alternative. This yields the striking conclusion that computationally efficient detection can be harder than computationally efficient estimation, revealing an unusual reverse detection–estimation gap: in a broad regime, computationally efficient estimation is possible at a smaller scale than computationally efficient detection. This phenomenon arises because the computational difficulty is driven not only by the statistical model, but also by the loss function itself: tensor spectral norm is NP-hard to compute. This feature makes the proposed open problems regarding computational lower bounds for estimation qualitatively different from the existing literature. Our results therefore uncover a new kind of computational–statistical gap.
1 Introduction
High-order moments and cumulants play a central role in modern statistics and learning theory. Beyond the covariance matrix, which captures only second-order dependence, higher-order moment and cumulant tensors encode rich non-Gaussian structure that underlies fundamental tasks such as independent component analysis (ICA) Comon (1994), latent variable estimation (Anandkumar et al., 2014), method-of-moments learning (Pearson, 1894), Gaussian mixture models (Wu and Yang, 2020), and multi-view representation learning (Anandkumar et al., 2012). In many of these problems, high-order tensors serve to identify hidden factors or to distinguish structured alternatives from Gaussian noise. Motivated by these developments, this paper initiates a systematic study of the statistical and computational limits of estimation and detection of high-order moment and cumulant tensors, without imposing additional structural assumptions such as sparsity or low rank.
Suppose we observe i.i.d. samples $X_{1},\ldots,X_{n}\in\mathbb{R}^{p}$ from an unknown distribution of a random vector $X$. We study the estimation and detection of the high-order moments and cumulants of $X$: for $d\geq 3$,
${\mathbfcal{M}}_{d}=\mathbb{E}[X^{\otimes d}],\qquad{\mathbfcal{K}}_{d}=\nabla^{d}\log\!\left(\mathbb{E}\,e^{t^{\top}X}\right)\big|_{t=0}.$
While ${\mathbfcal{M}}_{d}$ is the most direct extension of covariance to higher order, cumulants provide a more refined characterization of non-Gaussianity: in particular, all cumulants of order $d\geq 3$ vanish for Gaussian distributions, and cumulants are additive under independent summation. These properties make ${\mathbfcal{K}}_{d}$ a canonical object both in statistical theory and in computational pipelines that first whiten second-order structure and then extract higher-order deviations from Gaussianity.
A key feature of our setting is that we measure estimation error in the tensor spectral norm (operator norm),
$\|{\mathbfcal{T}}\|:=\sup_{u_{k}:\|u_{k}\|=1,k\in[d]}\Big\langle{\mathbfcal{T}},\ \bigotimes_{k\in[d]}u_{k}\Big\rangle,$
the natural analogue of matrix operator norm. This loss is the benchmark in covariance estimation because it uniformly controls directional quadratic forms and directly governs eigenvalue, eigenspace, and PCA perturbation. In the tensor setting, spectral norm likewise controls worst-case multilinear functionals, underlies perturbation guarantees for spectral and tensor power methods, and connects naturally to testing problems that distinguish Gaussian-like distributions from structured alternatives. By contrast, entrywise or Frobenius losses primarily average error over coordinates; for unstructured moment or cumulant tensors, they reduce essentially to high-dimensional mean estimation of the vectorized tensor and therefore miss the genuinely tensor-specific geometry captured by spectral norm.
When $d=2$, covariance estimation in spectral norm is well understood: the sample covariance achieves the optimal rate $\sqrt{p/n}$ under sub-Gaussian assumptions, while for heavy-tailed distributions, suitably truncated versions attain the same rate (Ke et al., 2019b; Koltchinskii and Lounici, 2017). For higher-order tensors, however, two striking differences arise.
  1. 1.
    Although the statistical minimax lower bound for estimating higher-order moment and cumulant tensors remains of order $\sqrt{p/n}\wedge 1$, the naive sample moment and sample cumulant estimators are no longer rate-optimal, even under sub-Gaussianity; see Theorems 2 and 4. Thus, unlike the covariance case, statistically optimal estimation requires more delicate procedures.
  2. 2.
    For $d\geq 3$, computing, certifying, or even approximating the tensor spectral norm is NP-hard in general (Hillar and Lim, 2013). This computational obstruction fundamentally changes the relationship between estimation and testing, and makes it impossible to directly transfer many matrix-based intuitions to the tensor setting.
These two facts suggest a basic question: what is the computationally optimal way to estimate high-order moments and cumulants in tensor spectral norm? To understand computational barriers for estimation, we study an associated detection problem. Our results show that this question exhibits a rich computational–statistical structure that is qualitatively different from both covariance estimation and standard planted-signal problems. In particular, we identify a regime in which computationally efficient detection is harder than computationally efficient estimation.
Overview of contributions.
We summarize the main contributions of the paper in the following informal theorem, followed by a more detailed discussion below.
Theorem 1 (Informal).
For fixed order $d\geq 3$, the problem of estimating and detecting high-order moment and cumulant tensors under tensor spectral norm exhibits the following behavior:
  1. 1.
    Statistical limit. Over broad classes of sub-Gaussian distributions, the minimax estimation error in tensor spectral norm satisfies
    $\inf_{\widehat{\mathbfcal{T}}}\sup_{{\mathbfcal{T}}}\mathbb{E}\|\widehat{\mathbfcal{T}}-{\mathbfcal{T}}\|\asymp\sqrt{p/n}\wedge 1.$
  2. 2.
    Computational limit for detection. For the associated testing problem, the low-degree threshold occurs at $n\asymp p^{d/2}$.
  3. 3.
    Reverse detection–estimation gap. There is a regime where computationally efficient detection is harder than computationally efficient estimation.
(i) Statistical limits. We establish minimax lower bounds for estimating both ${\mathbfcal{M}}_{d}$ and ${\mathbfcal{K}}_{d}$ in tensor spectral norm over broad classes of sub-Gaussian distributions. Specifically, the optimal statistical rate is
$\sqrt{p/n}\wedge 1,$
up to constants depending only on $d$; see Theorem 2. We then show that the sample moment and sample cumulant tensors generally fail to achieve this rate for $d\geq 3$: under $\beta$-subexponential tails, their operator-norm error scales as
$\max\Big\{\frac{p^{d/\beta}}{n},\ \sqrt{\frac{p}{n}}\Big\},$
which is sharp in its dependence on $(n,p)$ even in the sub-Gaussian case $\beta=2$; see Section 3.2. Moreover, we construct a statistically optimal (up to a logarithmic factor) estimator via a convex feasibility formulation over an $\varepsilon$-net of the sphere, showing that the minimax benchmark $\sqrt{p/n}$ is statistically attainable; see Section 3.3. Finally, we show that detection and consistent estimation have the same statistical limit. This establishes the statistical landscape of the problem.
(ii) A “reverse” detection–estimation gap. We next turn to the computational aspect of the problem. To understand when efficient procedures may or may not exist, we study an associated detection problem: given i.i.d. samples, test whether the $d$-th order cumulant tensor vanishes after whitening, versus whether it is bounded away from zero in tensor spectral norm. Using the low-degree polynomial framework, we show that detection is computationally hard in the regime
$n\ll p^{d/2},$
for a broad class of polynomial-time methods; see Section 4.2.
Moreover, our construction yields a family of testing problems for which low-degree methods fail even when the alternative is separated from the null by at least
$\sqrt{\frac{p^{d/2}}{n(\log(n))^{8}}}.$
On the other hand, the sample cumulant tensor is computationally efficient to compute, and under sub-Gaussian assumptions, its operator-norm error is of order
$\frac{p^{d/2}}{n},\qquad\text{when }p^{d/2}\ll n\ll p^{d-1}.$
Comparing these two scales reveals an unusual phenomenon: the error of an efficiently computable estimator is asymptotically smaller than the separation at which low-degree tests can reliably distinguish the null from the alternative. In this sense, computationally efficient detection can be strictly harder than computationally efficient estimation. This is in contrast to the more familiar detection–recovery gap in planted problems, where recovery is typically the harder task (Bresler and Jiang, 2023; Mao et al., 2023; Schramm and Wein, 2022).
The underlying reason is that an efficiently computable estimator does not automatically yield an efficiently computable test in our setting. Indeed, turning an estimator into a test requires certifying the size of the estimated tensor in spectral norm, and computing or approximating that norm is NP-hard in general. Thus the usual reduction from estimation to testing does not preserve computational efficiency here.
(iii) Open problems. Our results leave several basic questions open. Most notably, we conjecture that the low-degree threshold $n\asymp p^{d/2}$ marks the onset of efficient consistent estimation, and that in the intermediate regime $p^{d/2}\ll n\ll p^{d-1}$, the computationally optimal estimation error is strictly larger than the statistical minimax rate. If true, this would yield a nearly complete computational–statistical phase diagram for moment and cumulant tensor estimation in spectral norm.
What makes these questions unusual is that the main difficulty does not arise solely from the size of the parameter space or from the complexity of the underlying distribution class. Rather, it is built into the loss function itself. In many open problems on computationally constrained estimation, the target loss is easy to evaluate once an estimator is given, and the main challenge is to determine whether one can compute an estimator attaining the statistically optimal rate. Here, by contrast, the natural loss—the tensor spectral norm—is itself NP-hard to compute or even approximate in general. As a result, the standard reduction from estimation to testing is no longer algorithmically transparent: even if one can efficiently construct an estimator, it may still be computationally intractable to certify that the estimator is accurate in the relevant norm.
At a high level, our results reveal a new computational–statistical tension in high-order moment and cumulant tensor estimation under the stringent spectral-norm loss. They suggest the phase diagram in Figure 1.
  1. 1.
    When $n\ll p$, consistent estimation and detection are both statistically impossible.
  2. 2.
    When $p\ll n$, consistent estimation and detection are statistically possible, but the computational perspective depends further on the sample size:
    1. (a)
      When $p\ll n\ll p^{d/2}$, we prove that detection is computationally infeasible and conjecture that consistent estimation is computationally infeasible.
    2. (b)
      When $p^{d/2}\ll n\ll p^{d-1}$, there exist computationally efficient procedures, and we conjecture that the computationally optimal estimation error is ${p^{d/2}}/{n}$, which is strictly worse than the minimax statistical rate.
    3. (c)
      When $p^{d-1}\ll n$, the sample moment and sample cumulant tensors already achieve the statistically optimal rate $\sqrt{p/n}$.
Figure 1: Phase diagram of statistical and computational limits in moment and cumulant tensor estimation and detection
The rest of the paper is organized as follows. Section 1.1 reviews related work. Section 2 introduces notation and preliminaries. Section 3 establishes the statistical limits of moment and cumulant tensor estimation, including minimax lower bounds, upper bounds for sample moments and cumulants, and a statistically optimal estimator. Section 4 studies computational barriers via the low-degree framework, proves a detection lower bound, and develops polynomial-time testing procedures based on Frobenius-norm surrogates. Section 5 discusses the resulting open problems and conjectures, with particular emphasis on efficient consistent estimation and efficient rate-optimal estimation.
1.1 Related Work
We next discuss several lines of work related to this paper.
Independent component analysis (ICA)
In the ICA model, one observes a random vector of the form $X=AS$, where $A$ is an unknown deterministic mixing matrix and $S$ has entrywise independent coordinates. A key step in many ICA algorithms is to exploit fourth-order information, since after whitening, non-Gaussianity is often identified through deviations of the 4th moment (or 4th cumulant) from that of a Gaussian distribution. Consequently, both detection and estimation in ICA are intimately connected to the estimation of 4th moments and cumulants.
Recent works Auddy and Yuan (2025); Székely et al. (2024) establish sharp computational–statistical gaps for ICA: when $n<p$, reliable detection is statistically impossible, whereas for $n<p^{2}$ it becomes computationally hard under broad algorithmic frameworks. These thresholds are consistent with the phase transitions suggested by our general theory when specialized to $d=4$.
Despite this close connection, ICA fundamentally differs from the setting considered here, since the relevant 4th-order tensor in ICA exhibits strong algebraic structure induced by independence and the mixing model. In particular, the structured nature of the parameter often enables reductions from estimation to detection. In contrast, our work studies high-order moments and cumulants in a fully general (nonparametric) setting without imposing structural constraints such as independence, low rank, or sparsity. Technically, our low-degree lower bound in Section 4.2 can be viewed as a generalization of the low-degree polynomial (LDP) construction developed in Székely et al. (2024).
Tensor PCA
Tensor PCA concerns the estimation of a rank-one signal tensor embedded in noise. A canonical formulation assumes an observation model $Y=X+Z,$ where $X=a^{\otimes d}$ and $a\in\mathbb{R}^{p}$ is an unknown vector and $Z$ is a noise tensor (e.g., Gaussian). The computational hardness of tensor PCA has been extensively studied using a variety of frameworks, including reductions (Zhang and Xia, 2018), the low-degree polynomial framework (Kunisky et al., 2019b), memory- or communication-bounded models (Dudeja and Hsu, 2024), and the statistical query framework (Dudeja and Hsu, 2021a). These works collectively demonstrate the existence of a broad computationally hard regime for detection and estimation when the signal strength is below the algorithmic threshold, typically scaling as $\|X\|\ll p^{d/4}$ in the rank-one spiked tensor model.
Our setting differs in that we do not assume a spiked model or low-rank structure for the moment/cumulant tensor. Nevertheless, tensor PCA provides useful intuition: if one imposes a rank-one structure on the order-$d$ cumulant tensor (i.e., ${\mathbfcal{K}}_{d}\propto u^{\otimes d}$), then the tensor PCA detection threshold corresponds to $n\ll p^{d/2}$ in our sample-based formulation. As in ICA, the rank-one assumption enables a direct reduction from estimation to detection.
Robust covariance estimation
Robust covariance estimation under weak moment or heavy-tailed assumptions has been studied extensively. Catoni (2016) develops PAC-Bayesian and M-estimation type procedures that estimate the covariance operator via uniform control of quadratic forms. Wei and Minsker (2017) proposes computationally tractable covariance estimators for heavy-tailed distributions with strong nonasymptotic guarantees, complementing truncation- and median-of-means based approaches. Minsker (2018) constructs sub-Gaussian covariance estimators achieving sharp spectral-norm guarantees under weak tail assumptions, demonstrating how truncation and robust mean ideas yield optimal operator-norm performance in the matrix setting. Ke et al. (2019a) provides practically implementable, heavy-tail-robust covariance estimators with user-friendly tuning and nonasymptotic spectral-norm bounds. Mendelson and Zhivotovskiy (2020) establishes sharp robust covariance estimation under an $L_{4}$$L_{2}$ norm equivalence (finite-kurtosis type) condition.
Polynomial chaos and concentration
Beyond second-order structure, higher-order moment and cumulant tensors are closely tied to polynomial chaoses. The decoupling inequalities of Kwapien (1987) form a classical tool for reducing dependent multilinear forms to independent counterparts in high-degree chaos analysis. LATAŁA (2006) proves two-sided moment and tail bounds for Gaussian chaoses of arbitrary order, a foundational reference for modern polynomial and tensor concentration results. Adamczak and Wolff (2015) derives concentration inequalities for non-Lipschitz functions with bounded higher derivatives, yielding multilevel bounds for polynomials beyond the Lipschitz-concentration regime. Götze et al. (2021) establishes general concentration inequalities for polynomial functionals of independent random variables, providing tools directly relevant to high-degree moment and tensor analysis.
Moment methods
Moment-based methods are also central to statistical inference in Gaussian mixture models (GMMs). Wu and Yang (2020) gives a denoised method-of-moments procedure via projection to truncated moment space and proves statistical guarantees and adaptive optimality. Doss et al. (2020) highlights the role of moment tensors of the mixing distribution. Pereira et al. (2022) derives explicit formulas and discusses efficient implicit computations for GMM moment tensors. Khouja et al. (2022) proves that parameters of spherical Gaussian mixtures can be learned from empirical moments by forming symmetric moment tensors, and gives a linear-algebra-based decomposition algorithm. Kottler et al. (2025) studies how to recover Gaussian-mixture parameters in practice using the method of moments.
Computational hardness
From the algorithmic perspective, tensor norm computation and related inference problems are known to be computationally hard. Hillar and Lim (2013) shows that many basic tensor problems, including deciding and approximating the tensor spectral norm, are NP-hard. Hopkins et al. (2015) analyzes Tensor PCA through sum-of-squares (SoS) proofs, demonstrating nontrivial recovery and certification guarantees and characterizing SoS power in spiked tensor models. Hopkins et al. (2017) develops a general theory connecting SoS methods to spectral algorithms built from low-degree polynomial matrices for detecting planted structure, including tensor problems. Barak and Moitra (2016) applies SoS relaxations to noisy tensor completion, illustrating how SoS can certify polynomial objectives closely related to tensor norms. Potechin and Rajendran (2020) provides general machinery for proving SoS lower bounds and derives SoS lower bounds for Tensor PCA, complementing upper-bound results on random tensors. Wein (2023) studies average-case complexity for random tensor decomposition within the low-degree framework, identifying regimes where low-degree methods succeed or fail. Brennan et al. (2020) shows a close connection between statistical query algorithms and low-degree tests, supporting low-degree as a computational model for many high-dimensional testing problems. Dudeja and Hsu (2021b) establishes statistical query lower bounds for Tensor PCA, providing hardness evidence under another widely used restricted-algorithm framework. Hopkins and Steurer (2017b) develops a low-degree/SoS perspective for certain Bayesian inference tasks, including community detection. Luo and Zhang (2020) highlights hypergraphic planted clique detection as a central average-case hardness conjecture for tensor problems and raises the question of whether its hardness is computationally equivalent to that of classical planted clique detection.
2 Preliminaries and Notation
Notation
We use $c,C,C^{\prime}$, etc. to denote positive constants whose values may vary from line to line, and use subscripts to indicate their dependence on parameters. For example, $C_{\beta,d}$ denotes a constant that depends only on $\beta$ and $d$, and not on any other parameters such as $p$. For a positive integer $d$, write $d]=\{1,\ldots,d\}$. Vectors are denoted by lowercase letters, matrices by uppercase letters, and tensors by calligraphic letters such as ${\mathbfcal{T}}\in\mathbb{R}^{p_{1}\times\cdots\times p_{d}}$. For $x\in\mathbb{R}^{p}$, let $\|x\|$ be its Euclidean norm, and let $\mathbb{S}^{p-1}=\{x\in\mathbb{R}^{p}:\|x\|=1\}$ be the unit sphere. For any random variable or random vector $X$, $Y$, we write $X\overset{d}{\sim}Y$ if they have the same distribution. For a matrix $A$, let $\|A\|$ and $\|A\|_{F}$ denote its spectral norm and Frobenius norm, respectively. For a tensor ${\mathbfcal{T}}\in(\mathbb{R}^{p})^{\otimes d}$, its spectral norm is
$\|{\mathbfcal{T}}\|:=\sup_{u_{1},\ldots,u_{d}\in\mathbb{S}^{p-1}}\langle{\mathbfcal{T}},u_{1}\otimes\cdots\otimes u_{d}\rangle.$
We write $\|{\mathbfcal{T}}\|_{F}$ for the tensor Frobenius norm. For a random vector $X\in\mathbb{R}^{p}$ and an integer $r\geq 1$, let ${\mathbfcal{M}}_{r}=\mathbb{E}[X^{\otimes r}]$ denote the $r$-th moment tensor and let ${\mathbfcal{K}}_{r}$ denote the $r$-th cumulant tensor. For indices $i_{1},\ldots,i_{r}\in[p]$, the corresponding entries are written as $({\mathbfcal{M}}_{r})_{i_{1},\ldots,i_{r}}$ and $({\mathbfcal{K}}_{r})_{i_{1},\ldots,i_{r}}$.
For a real-valued random variable $Y$ and $\beta>0$, define the Orlicz norm by
$\|Y\|_{\Phi_{\beta}}:=\inf\{t>0:\mathbb{E}\exp(|Y|^{\beta}/t^{\beta})\leq 2\}.$
We call a random vector $X\in\mathbb{R}^{p}$ $\beta$-subexponential for some $\beta>0$ if
$\|v^{\top}X\|_{\Phi_{\beta}}:=\inf\Big\{t>0:\ \mathbb{E}\exp\big(|v^{\top}X|^{\beta}/t^{\beta}\big)\leq 2\Big\}\leq 1,\qquad\forall v\in\mathbb{S}^{p-1}.$
In particular, $\beta=2$ corresponds to the sub-Gaussian case and $\beta=1$ to the sub-exponential case.
Moment and cumulant tensors
We briefly review the definitions of moment and cumulant tensors and their relationship. Let $X=(X_{1},\ldots,X_{p})^{\top}\in\mathbb{R}^{p}$ be a random vector with finite moments up to order $d$. For $r\geq 1$, the $r$-th moment tensor is ${\mathbfcal{M}}_{r}=\mathbb{E}[X^{\otimes r}]\in(\mathbb{R}^{p})^{\otimes r}$, whose entries are $({\mathbfcal{M}}_{r})_{i_{1},\ldots,i_{r}}=\mathbb{E}[X_{i_{1}}\cdots X_{i_{r}}]$.
The cumulant tensors are defined from the cumulant generating function, assuming the cumulant generating function exists in a neighborhood of the origin: for $t\in\mathbb{R}^{p}$, $K_{X}(t):=\log\mathbb{E}\exp(t^{\top}X)$. The $r$-th cumulant tensor ${\mathbfcal{K}}_{r}\in(\mathbb{R}^{p})^{\otimes r}$ is defined entrywise by
$({\mathbfcal{K}}_{r})_{i_{1},\ldots,i_{r}}:=\frac{\partial^{r}K_{X}(t)}{\partial t_{i_{1}}\cdots\partial t_{i_{r}}}\Big|_{t=0}.$
In particular, ${\mathbfcal{K}}_{1}=\mathbb{E}[X]$, and when $X$ is centered, ${\mathbfcal{K}}_{2}$ and ${\mathbfcal{K}}_{3}$ coincide with the second and third central moment tensors.
To describe the general relation between moments and cumulants, we use set partitions. Let $\mathcal{P}([d])$ denote the collection of all partitions of $[d]$. A partition $\pi\in\mathcal{P}([d])$ is a collection of nonempty, pairwise disjoint subsets $\pi=\{B_{1},\ldots,B_{m}\}$, called the blocks of $\pi$, whose union is $[d]$. For a block $B=\{j_{1},\ldots,j_{r}\}\subset[d]$, we write $i_{B}:=(i_{j_{1}},\ldots,i_{j_{r}})$, where the indices are taken in increasing order. Then the moment–cumulant formula (Speed, 1983) states that
$({\mathbfcal{M}}_{d})_{i_{1},\ldots,i_{d}}=\sum_{\pi\in\mathcal{P}([d])}\prod_{B\in\pi}({\mathbfcal{K}}_{|B|})_{i_{B}}.$
Conversely, cumulants can be recovered from moments by Möbius inversion on the lattice of set partitions. Möbius inversion on the lattice of set partitions is a general combinatorial method for reversing a summation formula over all ways of partitioning a set, using coefficients given by the Möbius function of that partially ordered set. Here the partition lattice is ordered by refinement: $\pi\preceq\sigma$ if every block of $\pi$ is contained in a block of $\sigma$. The inversion formula (Speed, 1983) for this lattice yields
$({\mathbfcal{K}}_{d})_{i_{1},\ldots,i_{d}}=\sum_{\pi\in\mathcal{P}([d])}\mu(\pi)\prod_{B\in\pi}({\mathbfcal{M}}_{|B|})_{i_{B}},\qquad\mu(\pi)=(-1)^{|\pi|-1}(|\pi|-1)!,$
where $|\pi|$ denotes the number of blocks in $\pi$. Thus, cumulants are alternating polynomial combinations of lower-order moments, and moments are polynomial combinations of lower-order cumulants.
In particular, there exist tensor-valued polynomials $F_{d}$ and $G_{d}$, depending only on $d$, such that
${\mathbfcal{K}}_{d}={\mathbfcal{M}}_{d}+F_{d}({\mathbfcal{M}}_{1},\ldots,{\mathbfcal{M}}_{d-1}),\qquad{\mathbfcal{M}}_{d}={\mathbfcal{K}}_{d}+G_{d}({\mathbfcal{K}}_{1},\ldots,{\mathbfcal{K}}_{d-1}).$ (1)
That is, the difference between the $d$-th cumulant tensor and the $d$-th moment tensor depends only on tensors of orders strictly smaller than $d$. Throughout, products appearing in these polynomials are understood entrywise according to the partition formulas above.
3 Statistical Limit
3.1 Lower Bounds for Estimation and Detection
We begin by establishing the statistical minimax lower bounds for estimating high-order moment and cumulant tensors in spectral norm.
Theorem 2 (Lower bound for estimating $d$-th order cumulant and moment tensors).
Let $d\geq 3$ and $p>C_{1}$ for some absolute constant $C_{1}>0$. Define the class of sub-Gaussian random vectors as
$\mathcal{F}=\Big\{X\in\mathbb{R}^{p}:\|v^{\top}X\|_{\Phi_{2}}\leq 2,\ \forall v\in\mathbb{S}^{p-1}\Big\}.$
Then for any estimators based on $n$ i.i.d. samples, we have
$\inf_{\widehat{{\mathbfcal{K}}}}\sup_{X\in\mathcal{F}}\mathbb{E}\big\|\widehat{{\mathbfcal{K}}}-{\mathbfcal{K}}_{d}(X)\big\|\;\geq\;c_{d}\Big(\sqrt{\frac{p}{n}}\wedge 1\Big),$
and
$\inf_{\widehat{{\mathbfcal{M}}}}\sup_{X\in\mathcal{F}}\mathbb{E}\big\|\widehat{{\mathbfcal{M}}}-\mathbb{E}X^{\otimes d}\big\|\;\geq\;c_{d}\Big(\sqrt{\frac{p}{n}}\wedge 1\Big),$
where ${\mathbfcal{K}}_{d}(X)$ denotes the $d$-th order cumulant tensor of $X$ and $c_{d}>0$ is a constant depending only on $d$.
Next, we turn to the associated detection problem
$H_{0}:\ {\mathbfcal{T}}_{d}=0\qquad\text{vs.}\qquad H_{1}:\ \|{\mathbfcal{T}}_{d}\|\geq\kappa,$
where ${\mathbfcal{T}}_{d}$ denotes either the $d$-th moment tensor or the $d$-th cumulant tensor.
Theorem 2 shows that consistent estimation is impossible when $n\lesssim p$. We now show that the same scaling is also necessary for reliable testing: if $n\ll p$, then no test can distinguish the null from alternatives separated by less than order $\sqrt{p/n}\wedge 1$ in spectral norm.
Theorem 3 (Statistical lower bound for detection).
Let $d\geq 3$, and let
$\mathcal{F}=\Big\{X\in\mathbb{R}^{p}:\ \|v^{\top}X\|_{\Phi_{2}}\leq 2,\ \forall v\in\mathbb{S}^{p-1}\Big\}.$
For $\rho>0$, define the cumulant testing problem
$H_{0}^{(K)}:\ {\mathbfcal{K}}_{d}(X)=0\qquad\text{vs.}\qquad H_{1}^{(K)}(\rho):\ \|{\mathbfcal{K}}_{d}(X)\|\geq\rho,$
with $X\in\mathcal{F}$. Let
$\mathfrak{R}_{n}^{(K)}(\rho):=\inf_{\psi}\left\{\sup_{X\in\mathcal{F}:\,{\mathbfcal{K}}_{d}(X)=0}\mathbb{P}_{X}^{\otimes n}(\psi=1)+\sup_{X\in\mathcal{F}:\,\|{\mathbfcal{K}}_{d}(X)\|\geq\rho}\mathbb{P}_{X}^{\otimes n}(\psi=0)\right\},$
where the infimum is over all tests $\psi\in\{0,1\}$. Then there exist constants $c_{d},\eta_{d}>0$, depending only on $d$, such that
$\mathfrak{R}_{n}^{(K)}(\rho)\geq\eta_{d}\qquad\text{whenever}\qquad\rho\leq c_{d}\Big(\sqrt{p/n}\wedge 1\Big).$
Moreover, if $d$ is odd, the same statement holds for moment detection:
$H_{0}^{(M)}:\ \mathbb{E}X^{\otimes d}=0\qquad\text{vs.}\qquad H_{1}^{(M)}(\rho):\ \|\mathbb{E}X^{\otimes d}\|\geq\rho.$
That is, there exist constants $c_{d}^{\prime},\eta_{d}^{\prime}>0$, depending only on $d$, such that
$\mathfrak{R}_{n}^{(M)}(\rho)\geq\eta_{d}^{\prime}\qquad\text{whenever}\qquad\rho\leq c_{d}^{\prime}\Big(\sqrt{p/n}\wedge 1\Big).$
Remark 1 (Detection is trivial for even-order moment tensors.).
We restrict attention to odd-order moments throughout. If an even-order moment vanishes under $H_{0}$, then this necessarily implies $\mathbb{P}_{H_{0}}(X=0)=1$, rendering the testing problem degenerate. In particular, the lower bound is trivial in this case.
Thus, detection and consistent estimation share the same statistical lower bound in the present problem: up to constants, the critical statistical scale for both tasks is $\sqrt{p/n}\wedge 1$.
3.2 Sample Moment Tensors Can Be Suboptimal
We next study the most natural baseline estimator for moment and cumulant tensors in the absence of additional structural assumptions, namely the sample moment tensor. The goal of this subsection is to show that, unlike the covariance case, this estimator can be statistically suboptimal for higher-order estimation.
A first indication comes from the following lower bound. Proposition 3.1 in Al-Ghattas et al. (2025) shows that there exists a Gaussian random vector $X\in\mathbb{R}^{p}$ with sub-Gaussian norm 1 such that
$\mathbb{E}\Big\|\frac{1}{n}\sum_{i=1}^{n}X_{i}^{\otimes d}-\mathbb{E}[X^{\otimes d}]\Big\|\geq C_{d}\max\Big\{\frac{p^{d/2}}{n},\ \sqrt{\frac{p}{n}}\Big\}.$
Thus, even under sub-Gaussianity, the sample moment tensor can exhibit an operator-norm error strictly larger than the minimax rate when $d\geq 3$.
The next theorem shows that this lower bound is essentially sharp more generally under $\beta$-subexponential tails.
Theorem 4 (Deviation bound for sample moment tensors).
Let $X_{1},\ldots,X_{n}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}X\in\mathbb{R}^{p}$ be a $\beta$-subexponential random vector for some $\beta>0$. Then, with probability at least $1-\exp(-c_{\beta,d}p)$,
$\Big\|\frac{1}{n}\sum_{i=1}^{n}X_{i}^{\otimes d}-\mathbb{E}[X^{\otimes d}]\Big\|\leq C_{\beta,d}\max\Big\{\frac{p^{d/\beta}}{n},\ \sqrt{\frac{p}{n}}\Big\},$
and
$\Big\|\frac{1}{n}\sum_{i=1}^{n}(X_{i}-\bar{X})^{\otimes d}-\mathbb{E}[(X-\mu)^{\otimes d}]\Big\|\leq C_{\beta,d}\max\Big\{\frac{p^{d/\beta}}{n},\ \sqrt{\frac{p}{n}}\Big\},$
where $\bar{X}=n^{-1}\sum_{i=1}^{n}X_{i}$ and $\mu=\mathbb{E}X$.
Theorem 4 provides a high-probability operator-norm bound for sample moment tensors. In the covariance case $d=2$ and under sub-Gaussian tails $\beta=2$, the bound recovers the classical $\sqrt{p/n}$ rate, consistent with Koltchinskii and Lounici (2017). [^4] The effective rank appearing in Koltchinskii and Lounici (2017) is always bounded above by $p$, with equality attained in the worst case.
A key distinction between covariance estimation and higher-order moment tensor estimation emerges for $d\geq 3$. To compare with the minimax benchmark $\sqrt{p/n}\wedge 1$, consider the estimator
$\widehat{{\mathbfcal{T}}}=\begin{cases}\frac{1}{n}\sum_{i=1}^{n}X_{i}^{\otimes d},&\text{if }n\geq p,\\[3.00003pt] 0,&\text{if }n<p.\end{cases}$
When $d=2$, the inequality $p/n\lesssim\sqrt{p/n}$ holds whenever $n\geq p$, so this estimator already achieves the minimax-optimal $\sqrt{p/n}$ rate. In contrast, for $d\geq 3$, the additional term $p^{d/\beta}/n$ can dominate $\sqrt{p/n}$ over a wide range of sample sizes. Therefore, sample moment tensors are generally not statistically optimal for higher-order estimation.
For cumulant tensor estimation, we use the plug-in estimator induced by the moment–cumulant relation (1),
$\widehat{{\mathbfcal{K}}}_{d}=\widehat{{\mathbfcal{M}}}_{d}+F_{d}(\widehat{{\mathbfcal{M}}}_{1},\ldots,\widehat{{\mathbfcal{M}}}_{d-1}),$
where $\widehat{{\mathbfcal{M}}}_{k}=n^{-1}\sum_{i=1}^{n}X_{i}^{\otimes k}$ denotes the $k$-th order sample moment tensor. We refer to $\widehat{{\mathbfcal{K}}}_{d}$ as the sample cumulant tensor. Since Theorem 4 applies to each order $k\leq d$, it provides operator-norm bounds for all lower-order sample moment tensors appearing in the plug-in formula. Because $F_{d}$ is a fixed tensor polynomial in $({\mathbfcal{M}}_{1},\ldots,{\mathbfcal{M}}_{d-1})$, with degree depending only on $d$, these bounds together imply that $\|\widehat{{\mathbfcal{K}}}_{d}-{\mathbfcal{K}}_{d}\|$ obeys the same operator-norm rate, up to constants depending only on $d$.
3.3 Statistically Optimal Estimator
In the previous subsection, we showed that sample moments and cumulants fail to achieve the minimax-optimal rate for estimating higher-order ($d\geq 3$) moment and cumulant tensors. In this section, we construct an estimator whose operator-norm error matches the minimax lower bound established in Theorem 2, up to logarithmic factors.
Step 1.
We begin by estimating a fixed multilinear functional of the moment tensor. Fix unit vectors $u_{1},\ldots,u_{d}\in\mathbb{S}^{p-1}$ and consider the scalar
$Y:=\big\langle X^{\otimes d},\bigotimes_{j=1}^{d}u_{j}\big\rangle=\prod_{j=1}^{d}\langle X,u_{j}\rangle.$
If $X$ is $\beta$-subexponential, then $Y$ is $(\beta/d)$-subexponential. Consequently, when $\beta/d<1$ (e.g., $d\geq 3$ for sub-Gaussian vectors), achieving concentration with exponentially small failure probability requires truncation; see Lemma 6.
Theorem 5 (Estimation of a multilinear form).
Let $X_{1},\ldots,X_{n}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}X\in\mathbb{R}^{p}$ be a $\beta$-subexponential random vector for some $\beta>0$. Let ${\mathbfcal{M}}=\mathbb{E}X^{\otimes d}$ and fix $u_{1},\ldots,u_{d}\in\mathbb{S}^{p-1}$. For a sufficiently large constant $\zeta>0$, define
$\widetilde{\mathbfcal{M}}(u_{1},\ldots,u_{d})=\begin{cases}\displaystyle\frac{1}{n}\sum_{i=1}^{n}X_{i}^{\otimes d},&n>p^{2d/\beta-1},\\[6.00006pt] \displaystyle\frac{1}{n}\sum_{i=1}^{n}\bigotimes_{j=1}^{d}\big(X_{i}I_{\{|\langle X_{i},u_{j}\rangle|\leq R\}}\big),&p<n\leq p^{2d/\beta-1},\\[6.00006pt] 0,&n\leq p,\end{cases}$
where $R=\big(c_{\beta,\zeta}\log(n/p)\big)^{1/\beta}$. Then, with probability at least $1-C\exp(-\zeta p)$,
$\big|\langle\widetilde{\mathbfcal{M}}(u_{1},\ldots,u_{d})-{\mathbfcal{M}},\ \bigotimes_{j=1}^{d}u_{j}\rangle\big|\leq C_{\beta,d}^{*}\big(\sqrt{p/n}\wedge 1\big),$
where
$C_{\beta,d}^{*}=\begin{cases}C_{\beta,d},&n>p^{2d/\beta-1},\\ R^{d},&p<n\leq p^{2d/\beta-1},\\ C_{\beta,d},&n\leq p.\end{cases}$
Remark 2 (Heuristic for the $\sqrt{p/n}$ rate).
Estimating $Y$ is essentially a one-dimensional mean estimation problem. Heuristically, the optimal deviation bound satisfies
$\mathbb{P}\big(|\widehat{Y}-Y|\lesssim\sqrt{\log(1/\delta)/n}\big)\geq 1-\delta.$
Taking $\delta=\exp(-\Theta(p))$, as required for the union bound arguments below, yields the $\sqrt{p/n}$ rate in Theorem 5.
Step 2.
Next, we use multilinear forms to estimate the target tensor. Since the moment tensor ${\mathbfcal{M}}=\mathbb{E}X^{\otimes d}$ is symmetric, it suffices to control diagonal forms of the form $\langle{\mathbfcal{M}},u^{\otimes d}\rangle$. For $u\in\mathbb{S}^{p-1}$, define
$m_{u}:=\langle{\mathbfcal{M}},u^{\otimes d}\rangle,\qquad\widetilde{m}_{u}:=\big\langle\widetilde{\mathbfcal{M}}(u,\ldots,u),u^{\otimes d}\big\rangle.$
For each fixed $u\in\mathbb{S}^{p-1}$, Theorem 5 yields
$|m_{u}-\widetilde{m}_{u}|\leq\tau:=C_{\beta,d}^{*}(\sqrt{p/n}\wedge 1)$
with high probability.
Let $\mathcal{N}$ be an $\varepsilon$-net of $\mathbb{S}^{p-1}$ with $\varepsilon=1/(2d)$ and $|\mathcal{N}|\leq\exp(c_{d}p)$. Define the feasible set
$\mathcal{S}=\Big\{{\mathbfcal{A}}\in\mathbb{R}^{p^{d}}:|\langle{\mathbfcal{A}},u^{\otimes d}\rangle-\widetilde{m}_{u}|\leq\tau,\ \forall u\in\mathcal{N}\Big\}.$
Applying a union bound over $\mathcal{N}$, we obtain
$\mathbb{P}\Big(\sup_{u\in\mathcal{N}}|m_{u}-\widetilde{m}_{u}|\leq\tau\Big)\geq 1-C_{d}\exp(-C^{\prime}_{\beta,d}p),$
which implies
$\mathbb{E}X^{\otimes d}\in\mathcal{S},$
with probability at least $1-C\exp(-c^{\prime}_{\beta,d}p)$. In particular, $\mathcal{S}$ is nonempty on this event.
For any ${\mathbfcal{A}}\in\mathcal{S}$ and any $v\in\mathbb{S}^{p-1}$, choose $u\in\mathcal{N}$ such that $\|u-v\|\leq\varepsilon$. A standard Lipschitz argument yields
$|\langle{\mathbfcal{A}}-\mathbb{E}X^{\otimes d},v^{\otimes d}\rangle|\leq|\langle{\mathbfcal{A}}-\mathbb{E}X^{\otimes d},u^{\otimes d}\rangle|+d\varepsilon\|{\mathbfcal{A}}-\mathbb{E}X^{\otimes d}\|.$
Combining this with the definition of $\mathcal{S}$ and taking the supremum over $v\in\mathbb{S}^{p-1}$, we obtain
$\sup_{v\in\mathbb{S}^{p-1}}|\langle{\mathbfcal{A}}-\mathbb{E}X^{\otimes d},v^{\otimes d}\rangle|\leq 2\tau+\tfrac{1}{2}\|{\mathbfcal{A}}-\mathbb{E}X^{\otimes d}\|.$
Applying Lemma 5, we conclude that
$\|{\mathbfcal{A}}-\mathbb{E}X^{\otimes d}\|\leq 4\tau,$
which matches the minimax-optimal rate up to the logarithmic factor contained in $C_{\beta,d}^{*}$.
The construction above defines a feasible set of tensors satisfying the moment constraints on the $\varepsilon$-net, but does not specify a unique estimator. To obtain a canonical choice, we impose an additional convex objective. In particular, minimizing a convex functional over the feasible set $\mathcal{S}$ yields a well-defined convex optimization problem, while preserving the statistical rate established above.
Theorem 6 (Convex formulation for a statistically optimal estimator).
Suppose the conditions in Theorem 5 hold. Let
$\widehat{\mathbfcal{M}}\in\operatornamewithlimits{argmin}_{{\mathbfcal{A}}\in\mathcal{S}}\|{\mathbfcal{A}}\|_{F}^{2}.$
Then, with probability at least $1-C\exp(-c^{\prime}_{\beta,d}p)$,
$\|\widehat{\mathbfcal{M}}-\mathbb{E}X^{\otimes d}\|\leq C_{\beta,d}^{*}(\sqrt{p/n}\wedge 1),$
and for $p>C^{\prime}_{\beta,d}$,
$\mathbb{E}\|\widehat{\mathbfcal{M}}-\mathbb{E}X^{\otimes d}\|\leq C_{\beta,d}^{*}(\sqrt{p/n}\wedge 1).$
For cumulant estimation, let
$\widehat{{\mathbfcal{K}}}_{d}=\widehat{{\mathbfcal{M}}}_{d}+F_{d}(\widehat{{\mathbfcal{M}}}_{1},\ldots,\widehat{{\mathbfcal{M}}}_{d-1}),$
where $F_{d}$ is defined as in the moment–cumulant relation (1), and let ${\mathbfcal{K}}_{d}$ be the $d$-th cumulant tensor. Since $F_{d}$ is a fixed polynomial in the lower-order moment tensors, the same rate carries over to cumulant estimation. In particular, with probability at least $1-C\exp(-c^{\prime}_{\beta,d}p)$,
$\|\widehat{\mathbfcal{K}}_{d}-{\mathbfcal{K}}_{d}\|\leq C_{\beta,d}^{*}(\sqrt{p/n}\wedge 1),$
and for $p>C^{\prime}_{\beta,d}$,
$\mathbb{E}\|\widehat{\mathbfcal{K}}_{d}-{\mathbfcal{K}}_{d}\|\leq C_{\beta,d}^{*}(\sqrt{p/n}\wedge 1).$
3.4 Optimal Detection
We now turn from estimation to detection. In general, consistent estimation implies statistical detectability; see, for example, page 80 of Tsybakov (2009) and page 95 of Rigollet and Hütter (2023). Indeed, if an estimator $\hat{\mu}$ consistently estimates a scalar parameter $\mu$, then the test
$H_{0}:\ \mu=0\qquad\text{vs.}\qquad H_{1}:\ |\mu|>c$
can be carried out using the rule $\xi=\mathbb{I}_{\{|\hat{\mu}|>c/2\}}$. Consistency of $\hat{\mu}$ guarantees that both the type I and type II errors vanish asymptotically.
Applying this principle to our setting, the statistical upper bounds for estimation immediately imply detectability whenever $n\gg p$. In particular, the testing problem
$H_{0}:\ {\mathbfcal{T}}_{d}=0\qquad\text{vs.}\qquad H_{1}:\ \|{\mathbfcal{T}}_{d}\|\geq\kappa$
is statistically solvable (though it may be computationally intensive) in the regime $n\gg p$, where ${\mathbfcal{T}}_{d}$ denotes either the $d$-th moment tensor or the $d$-th cumulant tensor. Combined with the lower bounds in Theorems 2 and 3, this establishes that, up to constants, detection and consistent estimation have the same statistical threshold in our problem.
4 Computational Limit
In the previous section, we characterized the statistical limits of estimating $d$-th order moment and cumulant tensors. In particular, for $\beta$-subexponential random vectors, we showed that the minimax-optimal estimation error in tensor spectral norm scales as $\sqrt{p/n}\wedge 1$. Achieving this optimal rate, however, requires solving a convex optimization problem with exponentially many constraints, which renders the resulting estimator computationally infeasible in high-dimensional settings. In the following sections, we investigate the computational limits of moment and cumulant tensor detection and estimation.
4.1 Low-degree Polynomial Framework
To study computational complexity, we adopt the low-degree polynomial (LDP) framework. Originating from the analysis of the sum-of-squares (SoS) hierarchy (Hopkins et al., 2017; Hopkins and Steurer, 2017a; Barak et al., 2019), the LDP framework has become a central tool for characterizing computational barriers in high-dimensional statistics. It has been successfully applied to a wide range of problems, including planted clique detection (Barak et al., 2019), tensor PCA (Hopkins, 2018), tensor regression (Luo and Zhang, 2024), among many others.
Consider a hypothesis testing problem with null distribution $\mathbb{Q}$ and alternative distribution $\mathbb{P}$. Under standard regularity conditions, the likelihood ratio test is asymptotically optimal by the Neyman–Pearson lemma. The LDP framework studies the projection of the likelihood ratio $L:=\frac{d\mathbb{P}}{d\mathbb{Q}}$ onto the space of polynomials of degree at most $D$. Specifically, it considers the norm
$\|L_{\leq D}\|:=\|\mathcal{P}_{\leq D}L\|,$
where $\mathcal{P}_{\leq D}$ denotes the orthogonal projection onto degree-$D$ polynomials and the norm is defined by $\|f\|:=(\mathbb{E}_{X\sim\mathbb{Q}}[f(X)^{2}])^{1/2}$.
Intuitively, $L_{\leq D}$ captures the optimal degree-$D$ polynomial statistic for distinguishing $\mathbb{P}$ from $\mathbb{Q}$, and the quantity $\|L_{\leq D}\|$ measures the extent to which low-degree polynomials can separate the two distributions. When $\|L_{\leq D}\|$ remains asymptotically bounded as the dimension grows, no degree-$D$ polynomial test can reliably distinguish $\mathbb{P}$ from $\mathbb{Q}$.
A broad class of algorithmic paradigms—including spectral methods, approximate message passing, statistical query algorithms, and other polynomial-time procedures—have been shown to be captured by low-degree polynomials. Consequently, the LDP framework provides a rigorous characterization of the computational limits of these algorithmic classes (Wein, 2025). Recent counterexamples show that bounded low-degree likelihood ratio norms do not, by themselves, rule out all efficient algorithms in complete generality (Buhai et al., 2025). Nevertheless, the low-degree framework continues to serve as a robust organizing principle for understanding statistical-computational tradeoffs across a broad range of structured inference problems (Holmgren and Wein, 2020; Buhai et al., 2025; Wein, 2025). We refer the interested reader to the surveys (Kunisky et al., 2019a; Wein, 2025) for comprehensive treatments of the LDP framework and its implications.
4.2 Detection Lower Bound
In this subsection, we apply the LDP framework described in Section 4.1 to study the computational hardness of the following detection problem: given i.i.d. samples from a random vector, determine whether its $d$-th order cumulant tensor is zero.
To this end, we construct a family of distributions whose lower-order cumulants match those of a Gaussian distribution, while exhibiting a nontrivial $d$-th order cumulant. This construction ensures that any algorithm must exploit genuinely high-order information to distinguish the alternative from the null. Specifically, let
$X_{i}=\sqrt{a/p}W_{i}U+Z_{i},\quad i\in[n],$
where $Z_{i}\overset{iid}{\sim}\mathcal{N}(0,I_{p})$, $W_{i}\in\mathbb{R}$ are i.i.d. random variables drawn from the distribution constructed in Lemma 3, $U\in\mathbb{R}^{p}$ is a fixed vector satisfying $\|U\|=\sqrt{p}$ (sampled from some prior), and $a>0$ is a constant to be specified later. Conditioning on $U$, the cumulants of $X_{i}$ satisfy
${\mathbfcal{K}}_{1}(X)=0;\;{\mathbfcal{K}}_{2}(X)=I_{p}+\frac{a}{p}UU^{\top};\;$
${\mathbfcal{K}}_{m}(X)=0,\ \text{for $m\in[3,d-1]$};\;{\mathbfcal{K}}_{d}(X)=c_{d}\left(\frac{a}{p}\right)^{d/2}U^{\otimes d},$
where $c_{d}$ is a constant depending only on $d$. Let
$S=I_{p}-\frac{a}{1+a+\sqrt{1+a}}\frac{UU^{\top}}{p}$
be the whitening matrix for ${\mathbfcal{K}}_{2}(X)$. Conditioning on $U$, the first d1d-1 cumulants of the transformed random vector $SX$ coincide with those of a standard Gaussian distribution, while
${\mathbfcal{K}}_{d}(SX)=c_{d}\left(\frac{a}{p}\right)^{d/2}(SU)^{\otimes d}.$
A direct calculation shows that
$\|SU\|^{2}=\frac{p}{1+a},$
and therefore
$\|{\mathbfcal{K}}_{d}(SX)\|=c_{d}\left(\frac{a\|SU\|^{2}}{p}\right)^{d/2}=c_{d}\left(\frac{a}{1+a}\right)^{d/2}.$
Thus, for any fixed constant $a>0$ and $\|U\|=\sqrt{p}$, the spectral norm of the order-$d$ cumulant of $SX$ is bounded below by a positive constant $C_{d}>0$ depending only on $d$.
We now formulate the following hypothesis testing problem based on $n$ i.i.d. samples:
$H_{0}:Y_{i}\overset{d}{\sim}Z_{i},\quad\text{vs. }\quad H_{1}:Y_{i}\overset{d}{\sim}SX_{i}.$ (2)
The low-degree likelihood ratio for this testing problem is controlled by the following lemma.
Lemma 1 (LDP norm upper bound).
$\|L_{\leq D}\|^{2}\leq\sum_{0\leq m\leq D}\left(\frac{a}{1+a}\frac{C_{d}m^{4}n^{1/d}}{p^{1/2}}\right)^{m}.$
As a consequence, we have the following computational lower bound.
Theorem 7 (Low-degree hardness for detecting a nonzero $d$-th order cumulant).
Fix $d\geq 3$ and consider the testing problem (2). Then there exists a constant $c_{d}>0$, depending only on $d$, such that for any $\varepsilon>0$, if
$p\geq n^{2/d+\varepsilon},\qquad D\leq c_{d}\,n^{\varepsilon/8},$
then the degree-$D$ low-degree likelihood ratio satisfies
$\|L_{\leq D}\|=O(1).$
Consequently, there is no sequence of degree-$D$ polynomial tests whose sum of Type I and Type II errors converges to zero.
4.3 A “Reverse” Detection–Estimation Gap
A useful distinction in this area is between detection and estimation (or recovery). In many planted or signal-plus-noise models, these two tasks can exhibit different computational thresholds. Recent work in the low-degree framework has made this distinction precise. In particular, Schramm and Wein (2022) developed general low-degree lower bounds for estimation and recovery, and showed that for problems such as planted submatrix and planted dense subgraph, estimation can be computationally hard even when detection is easy. Luo and Gao (2024) identified a setting in which subtensor detection is easier than estimation. See also Mao et al. (2023) for an explicit detection–estimation gap in planted dense cycles, and Bresler and Jiang (2023) for reduction-based evidence of such a gap in planted dense subgraph. By contrast, in some models the two thresholds coincide; for example, in the spiked Wigner model, detection and estimation undergo the same phase transition (El Alaoui et al., 2020).
In the above examples, estimation is shown to be harder than detection, which is consistent with the intuition that recovering a signal should require more information than merely detecting its presence. The phenomenon described below goes in the opposite direction. We identify a regime in which there exists a computationally efficient estimator whose error is asymptotically smaller than the separation at which low-degree detection appears possible. In this sense, computationally efficient detection becomes strictly harder than computationally efficient estimation.
Now consider the regime $p<n<p^{d-1}$. In Lemma 1, if we choose
$a=\frac{p^{1/2}}{n^{1/d}(\log(n))^{8}},$
then $\|L_{\leq D}\|^{2}=O_{n}(1)$ with $D\asymp(\log(n))^{2}$. Meanwhile,
$\|{\mathbfcal{K}}_{d}(SX)\|=c_{d}\left(\frac{a}{1+a}\right)^{d/2}\gtrsim\sqrt{\frac{p^{d/2}}{n(\log(n))^{8}}}:=D_{\operatorname{det}}.$
Thus low-degree tests fail to distinguish $H_{0}$ and $H_{1}$ even when the alternative is separated from the null by at least $D_{\operatorname{det}}$ in tensor spectral norm. Under the usual low-degree conjecture, this provides evidence that no polynomial-time test can distinguish $H_{0}$ and $H_{1}$ at that scale.
On the other hand, Theorem 4 shows that the empirical cumulant tensor satisfies the error bound
$\|\widehat{{\mathbfcal{K}}}_{d}-{\mathbfcal{K}}_{d}\|\lesssim\frac{p^{d/2}}{n}:=D_{\operatorname{est}}$
whenever $n<p^{d-1}$. In particular, if $n=p^{d/2+\varepsilon}$ for some $\varepsilon\in(0,d/2-1)$, then
$D_{\operatorname{est}}\ll D_{\operatorname{det}}.$
Therefore, there exists a computationally efficient estimator whose estimation error is much smaller than the null–alternative separation required for computationally efficient detection. Consequently, in this regime (where both $D_{\operatorname{est}}$ and $D_{\operatorname{det}}$ vanish asymptotically), computationally efficient detection is strictly harder than computationally efficient estimation.
This may initially appear counterintuitive. The key point is that we are comparing computationally efficient procedures. Although the estimator $\widehat{{\mathbfcal{K}}}_{d}$ can be computed efficiently, converting it into an efficient detector is not straightforward. A natural decision rule would require certifying whether $\|\widehat{{\mathbfcal{K}}}_{d}\|$ exceeds a threshold, or more generally whether $\widehat{{\mathbfcal{K}}}_{d}$ is close to the true tensor in tensor spectral norm. However, these certification steps involve tensor spectral norms, whose exact computation is NP-hard in general (Hillar and Lim, 2013). Thus an efficiently computable estimator does not automatically yield an efficiently computable test. See Section 5 for further discussion.
Finally, if computational constraints are removed, estimation remains at least as hard as detection, as discussed in Section 3.4: any estimator with sufficiently small error can be converted into a test (see, for example, page 80 of Tsybakov (2009) and page 95 of Rigollet and Hütter (2023)). The point here is precisely that such a reduction need not preserve computational efficiency.
4.4 Detection Methods
Moment tensor.
Next, we consider a computationally efficient test for the moment tensor detection problem
$H_{0}:\ {\mathbfcal{M}}_{d}=\mathbb{E}[X^{\otimes d}]=0\qquad\text{vs.}\qquad H_{1}:\ \|{\mathbfcal{M}}_{d}\|\geq\kappa,$ (3)
based on i.i.d. samples $X_{1},\ldots,X_{n}$, where $\kappa>0$ is a fixed constant. Our goal is to construct a test statistic that is computable in polynomial time and achieves consistency when $n\gg p^{d/2}$. We restrict attention to odd orders $d$, since for even $d$, the condition ${\mathbfcal{M}}_{d}=0$ implies $X=0$, and the testing problem becomes trivial.
To make the test computationally efficient, we need to avoid evaluating the tensor spectral norm. Note that
$\|{\mathbfcal{M}}_{d}\|_{F}\geq\|{\mathbfcal{M}}_{d}\|.$
Therefore, any procedure that detects whether $\|{\mathbfcal{M}}_{d}\|_{F}$ is bounded away from zero can also distinguish $H_{1}$ from $H_{0}$. This suggests targeting the Frobenius norm instead.
However, if one first estimates the full tensor ${\mathbfcal{M}}_{d}$ and then computes its Frobenius norm, the resulting estimation error is of order $\sqrt{p^{d}/n}$, which yields consistency only when $n\gg p^{d}$ rather than when $n\gg p^{d/2}$. The key idea is therefore not to estimate ${\mathbfcal{M}}_{d}$ itself and then take its Frobenius norm, but instead to estimate $\|{\mathbfcal{M}}_{d}\|_{F}$ directly.
Indeed,
$\mathbb{E}\big[(X_{1}^{\top}X_{2})^{d}\big]$ $=\mathbb{E}\Big[\Big(\sum_{j=1}^{p}(X_{1})_{j}(X_{2})_{j}\Big)^{d}\Big]$
$=\sum_{i_{1},\ldots,i_{d}\in[p]}\mathbb{E}\big[(X_{1})_{i_{1}}\cdots(X_{1})_{i_{d}}\big]^{2}=\|{\mathbfcal{M}}_{d}\|_{F}^{2}.$
Motivated by this identity, define the test statistic
$V_{n}:=\frac{1}{n(n-1)}\sum_{i\neq j}(X_{i}^{\top}X_{j})^{d}.$ (4)
Then $\mathbb{E}V_{n}=\|{\mathbfcal{M}}_{d}\|_{F}^{2}$. Moreover, this statistic is computationally efficient, with total arithmetic cost $O(n^{2}p)$. This leads to the following polynomial-time consistent detection procedure.
Theorem 8 (Polynomial-time consistent test for vanishing $d$-th moment).
Suppose $d$ is odd. Let $X_{1},\ldots,X_{n}$ be i.i.d. copies of a random vector $X\in\mathbb{R}^{p}$ with sub-Gaussian norm $K$ for some constant $K>0$. Consider the hypothesis test (3) with test statistic $V_{n}$ defined in (4). Assume in addition that every alternative under $H_{1}$ satisfies
$\text{Var}\big(\langle{\mathbfcal{M}}_{d},X^{\otimes d}\rangle\big)\leq C_{1}p^{d/2},$ (5)
for some constant $C_{1}>0$. Let
$\tau_{n}=A\frac{p^{d/2}\log n}{n},$
where $A>0$ is a sufficiently large constant, and define the decision rule
$\phi_{n}:=\mathbf{1}_{\{V_{n}>\tau_{n}\}}.$
Then for fixed $d$:
  1. 1.
    $\phi_{n}$ is computable in $O(n^{2}p)$ arithmetic operations.
  2. 2.
    Under $H_{0}$,
    $\mathbb{P}_{H_{0}}(\phi_{n}=1)\to 0\qquad\text{whenever}\qquad\frac{p^{d/2}\log n}{n}\to 0.$
  3. 3.
    Under $H_{1}$,
    $\mathbb{P}_{H_{1}}(\phi_{n}=0)\to 0\qquad\text{whenever}\qquad\frac{p^{d/2}\log n}{n}\to 0.$
In particular, under (5), the test is polynomial-time computable and consistent whenever
$n\gg p^{d/2}\log n.$
Remark 3.
The additional condition $\text{Var}\big(\langle{\mathbfcal{M}}_{d},X^{\otimes d}\rangle\big)\leq C_{1}p^{d/2}$ is imposed to control the non-degenerate part of the U-statistic under $H_{1}$. For many structured alternatives, this condition is mild. For example, if ${\mathbfcal{M}}_{d}=\lambda\,u^{\otimes d}$ with $|\lambda|=O(1)$ and $u^{\top}X$ having bounded $2d$-th moment, then
$\langle{\mathbfcal{M}}_{d},X^{\otimes d}\rangle=\lambda(u^{\top}X)^{d},\quad\text{and}\quad\text{Var}\big(\langle{\mathbfcal{M}}_{d},X^{\otimes d}\rangle\big)=O(1).$
More generally, the assumption is natural for low-rank or spiked alternatives, including rank-one alternatives of the type used in our computational lower-bound construction.
Cumulant tensor.
We now consider the analogous detection problem for cumulant tensors. Suppose the observations are centered and whitened, and let $Y_{1},\ldots,Y_{n}$ be i.i.d. copies of a random vector $Y\in\mathbb{R}^{p}$ satisfying $\mathbb{E}Y=0$ and $\mathbb{E}[YY^{\top}]=I_{p}$. Let
${\mathbfcal{K}}_{r}\in(\mathbb{R}^{p})^{\otimes r},\qquad r=1,\ldots,d,$
denote the cumulant tensors of $Y$, and suppose that the lower-order cumulants ${\mathbfcal{K}}_{1},\ldots,{\mathbfcal{K}}_{d-1}$ are known. We are interested in testing
$H_{0}:\ {\mathbfcal{K}}_{d}=0\qquad\text{vs.}\qquad H_{1}:\ \|{\mathbfcal{K}}_{d}\|\geq\kappa,$ (6)
where $\kappa>0$ is a fixed constant.
In the moment case, the basic unbiased feature is $X^{\otimes d}$. For cumulants, the analogous object is a cumulant-corrected polynomial feature map. Since the lower-order cumulants are assumed known, we may form a known degree-$d$ tensor-valued polynomial feature map
$\Psi_{d}(\,\cdot\,;{\mathbfcal{K}}_{1},\ldots,{\mathbfcal{K}}_{d-1}):\mathbb{R}^{p}\to(\mathbb{R}^{p})^{\otimes d},$
constructed from the multivariate moment–cumulant relation, such that
$\mathbb{E}[\Psi_{d}(Y;{\mathbfcal{K}}_{1},\ldots,{\mathbfcal{K}}_{d-1})]={\mathbfcal{K}}_{d}.$ (7)
To simplify notation, we write
$\Psi_{d}(y):=\Psi_{d}(y;{\mathbfcal{K}}_{1},\ldots,{\mathbfcal{K}}_{d-1}).$
For example, when $d=3$, since $Y$ is centered and whitened, one may take
$\Psi_{3}(y)=y^{\otimes 3}-\operatorname{sym}(y\otimes I_{p}),$
so that $\mathbb{E}[\Psi_{3}(Y)]={\mathbfcal{K}}_{3}$. For general fixed $d$, $\Psi_{d}$ can be defined recursively using the known lower-order cumulants.
Define the U-statistic
$V_{n}^{(\mathrm{cumu})}:=\frac{1}{n(n-1)}\sum_{i\neq j}\langle\Psi_{d}(Y_{i}),\Psi_{d}(Y_{j})\rangle.$ (8)
Then, by independence,
$\mathbb{E}V_{n}^{(\mathrm{cumu})}=\Big\langle\mathbb{E}[\Psi_{d}(Y_{1})],\mathbb{E}[\Psi_{d}(Y_{2})]\Big\rangle=\|{\mathbfcal{K}}_{d}\|_{F}^{2}.$
Hence $V_{n}^{(\mathrm{cumu})}$ directly estimates $\|{\mathbfcal{K}}_{d}\|_{F}^{2}$, and therefore can be used to test (6).
Theorem 9 (Polynomial-time consistent test for vanishing $d$-th order cumulant tensor).
Let $Y_{1},\ldots,Y_{n}$ be i.i.d. centered and whitened random vectors in $\mathbb{R}^{p}$, and let ${\mathbfcal{K}}_{d}$ denote the $d$-th cumulant tensor of $Y$. Assume that the lower-order cumulants ${\mathbfcal{K}}_{1},\ldots,{\mathbfcal{K}}_{d-1}$ are known, and let $\Psi_{d}$ be a known degree-$d$ tensor-valued polynomial map satisfying (7). Consider the hypothesis test (6) with test statistic $V_{n}^{(\mathrm{cumu})}$ defined in (8).
Assume in addition that
$\mathbb{E}\Big[\langle\Psi_{d}(Y_{1}),\Psi_{d}(Y_{2})\rangle^{2}\Big]\leq C_{0}p^{d}$ (9)
for some constant $C_{0}>0$, and that every alternative under $H_{1}$ satisfies
$\text{Var}\big(\langle{\mathbfcal{K}}_{d},\Psi_{d}(Y)\rangle\big)\leq C_{1}p^{d/2}$ (10)
for some constant $C_{1}>0$. Let
$\tau_{n}=A\frac{p^{d/2}\log n}{n},$
where $A>0$ is a sufficiently large constant, and define
$\phi_{n}^{(\mathrm{cumu})}:=\mathbf{1}_{\{V_{n}^{(\mathrm{cumu})}>\tau_{n}\}}.$
Then for fixed $d$:
  1. 1.
    $\phi_{n}^{(\mathrm{cumu})}$ is computable in polynomial time in $(n,p)$. More precisely, if each evaluation of $\Psi_{d}(y)$ requires at most $O(p^{d})$ arithmetic operations, then $V_{n}^{(\mathrm{cumu})}$ can be computed in $O(n^{2}p^{d})$ arithmetic operations.
  2. 2.
    Under $H_{0}$,
    $\mathbb{P}_{H_{0}}\big(\phi_{n}^{(\mathrm{cumu})}=1\big)\to 0\qquad\text{whenever}\qquad\frac{p^{d/2}\log n}{n}\to 0.$
  3. 3.
    Under $H_{1}$,
    $\mathbb{P}_{H_{1}}\big(\phi_{n}^{(\mathrm{cumu})}=0\big)\to 0\qquad\text{whenever}\qquad\frac{p^{d/2}\log n}{n}\to 0.$
In particular, under (9) and (10), the test is polynomial-time computable and consistent whenever
$n\gg p^{d/2}\log n.$
Remark 4.
The assumption $\text{Var}\big(\langle{\mathbfcal{K}}_{d},\Psi_{d}(Y)\rangle\big)\leq C_{1}p^{d/2}$ is the cumulant analogue of (5). It controls the non-degenerate part of the U-statistic under $H_{1}$ and is natural for many structured alternatives, such as low-rank or spiked cumulant tensors.
The second assumption $\mathbb{E}[\langle\Psi_{d}(Y_{1}),\Psi_{d}(Y_{2})\rangle^{2}]\leq C_{0}p^{d}$ is a second-moment bound. It holds in many sub-Gaussian settings because $\Psi_{d}(Y)$ is a fixed degree-$d$ polynomial in $Y$ whose coefficients depend only on the known lower-order cumulants.
Remark 5.
The feature map $\Psi_{d}$ may be interpreted as a cumulant-corrected version of $Y^{\otimes d}$, obtained by removing the lower-order contributions in the moment–cumulant formula. When the lower-order cumulants are known, this map is fully determined and hence explicitly available to the testing procedure.
More generally, the theorem does not depend on this particular construction: it remains valid for any known feature map $\Psi_{d}$ such that $\mathbb{E}[\Psi_{d}(Y)]={\mathbfcal{K}}_{d}$. Thus, knowledge of the lower-order cumulants is a convenient sufficient condition for constructing $\Psi_{d}$, but not a necessary condition for the validity of the result.
5 Open Problems
The previous section characterized the computational limits of the detection problem. A natural next question is to understand the computational limits of estimation, and in particular to establish computational lower bounds. Compared with detection, however, this question is substantially more subtle. Existing approaches from the literature do not directly apply here, because estimation under tensor spectral norm has several features that differ qualitatively from the standard settings in which computational lower bounds are well understood. As a result, the computational complexity of estimation in our problem remains largely open.
5.1 Consistent Estimation
In the literature, the computational hardness of estimation problems is often established by reduction to the corresponding hypothesis testing problem. This reduction can be made formal (see, e.g., Section 17 of Brennan and Bresler (2020)), but the underlying logic is straightforward and the same as the one discussed in Section 3.4: if an estimator $\hat{\mu}$ consistently estimates a scalar parameter $\mu$, then one can test $H_{0}:\mu=0$ vs. $H_{1}:\mu>c$ using the decision rule $\xi=\mathbb{I}_{\{|\hat{\mu}|>c/2\}}$. Consistency of $\hat{\mu}$ guarantees that $\xi$ converges in probability to the correct decision, implying that the computational complexity of testing provides a lower bound for that of estimation.
This reasoning, however, does not directly apply to moment or cumulant tensor estimation. The key obstruction is that the natural loss function of interest—the tensor spectral norm—is itself NP-hard to compute or even approximate. Consequently, even if a consistent estimator $\widehat{{\mathbfcal{T}}}$ of a tensor parameter $\mathbfcal{T}$} can be computed efficiently, the corresponding decision rule $\xi=I_{\{\|\widehat{\mathbfcal{T}}\|>c/2\}}$ may remain computationally intractable.
Despite this difficulty, there is a striking alignment between the computational threshold for hypothesis testing established in Section 4.2 and the upper bound in Theorem 4. Specifically, when $n\gg p^{d/2}$, the sample cumulant tensor provides a computationally efficient consistent estimator, whereas when $n\ll p^{d/2}$, hypothesis testing becomes computationally hard. This motivates the following conjecture.
Conjecture 1.
When $n\ll p^{d/2}$, there is no efficient algorithm that can consistently estimate the order-$d$ cumulant or moment tensors of any sub-Gaussian random vector.
Proving Conjecture 1 is challenging, primarily due to the tensor spectral norm. To the best of our knowledge, very few works directly study computational lower bounds for estimation problems involving losses that are themselves hard to compute, making average-case reductions difficult to construct. Moreover, existing frameworks for establishing computational hardness of estimation—such as the low-degree polynomial or statistical query frameworks—typically require the loss to decompose into a sum of coordinate-wise losses (e.g., the Frobenius norm), and therefore do not apply here. Additional obstacles arise from the fact that the parameter space has dimension $p^{d}$, which complicates the application of memory- or communication-constrained models.
If Conjecture 1 holds, then combined with the results of Section 3.2, it yields a complete computational-statistical phase diagram for consistent estimation.
  • When $n\ll p$, the problem is statistically impossible.
  • When $p\ll n\ll p^{d/2}$, consistent estimation is statistically possible but computationally intractable.
  • When $n\gg p^{d/2}$, the sample moment (or cumulant) tensor is a computationally efficient consistent estimator.
One piece of indirect evidence supporting Conjecture 1 is that, if one attempts to adapt the spectrally truncated moment estimator, which is consistent for heavy-tailed covariance estimation whenever $n\gg p$, then a naive bias–variance decomposition for higher-order moments fails to deliver the desired rate. Specifically, the truncation bias and the stochastic fluctuation cannot be controlled simultaneously at the target scale.
5.2 Optimal Estimation
Recall from Section 3 that the minimax-optimal error rate for estimating $d$-th order moment tensors in spectral norm is $\sqrt{p/n}$. In contrast, when $p^{d/2}\ll n\ll p^{d-1}$, the sample moment tensor estimator attains an error rate of $p^{d/2}/n$, which is strictly suboptimal. In many statistical problems, once consistent estimation becomes possible, one can further refine the estimator to achieve the optimal rate—often by using a consistent estimator as initialization for an iterative refinement procedure. Examples of this paradigm appear in tensor problems with low-rank structure (Zhang and Xia, 2018; Tang et al., 2025).
In the moment and cumulant estimation setting, however, we do not impose any additional structural constraints such as low rankness, which significantly limits opportunities for algorithmic refinement beyond the sample moment tensors. A related line of work concerns covariance estimation in spectral norm under heavy-tailed distributions, where suitably truncated sample covariance estimators are known to be optimal (Minsker and Wei, 2018; Ke et al., 2019b). These techniques rely crucially on spectral properties of matrices and do not extend to higher-order tensors, highlighting a fundamental gap in current methodology.
On the other hand, there are notable exceptions where a parameter can be consistently estimated but cannot be estimated at the optimal rate by any efficient algorithm (Luo and Gao, 2024; Zhang et al., 2014). This leads us to the following conjecture.
Conjecture 2.
When $p^{d/2}\ll n\ll p^{d-1}$, the computationally optimal error rate for estimating $d$-th order moment or cumulant tensors in spectral norm is $p^{d/2}/n$.
The main obstacle to resolving Conjecture 2 is the same as for Conjecture 1: the lack of general frameworks capable of establishing computational lower bounds for estimation problems with hard-to-compute losses.
6 Discussion
In this paper, we studied the estimation and detection of high-order moment and cumulant tensors under tensor spectral norm, and uncovered a computational–statistical landscape that is qualitatively different from the matrix case and from more classical planted problems. On the statistical side, we showed that the minimax-optimal estimation rate is $\sqrt{p/n}\wedge 1$, while the naive sample moment and cumulant tensors are generally suboptimal for $d\geq 3$. On the computational side, low-degree evidence suggests that detection is hard when $n\ll p^{d/2}$, and this leads to an unusual reverse detection–estimation gap.
Our results leave open the computational complexity of estimation itself. In many high-dimensional problems, computational lower bounds for estimation are derived through reductions from testing. In the present setting, however, such reductions are no longer algorithmically transparent, because the relevant loss function is the tensor spectral norm, which is NP-hard to compute or approximate in general. As a result, even if an estimator can be computed efficiently, certifying that it achieves small error in the target norm may itself be computationally intractable.
This suggests several directions for future work. A first goal is to determine whether the threshold $n\asymp p^{d/2}$ indeed marks the onset of computationally efficient consistent estimation. A second goal is to characterize the computationally optimal estimation error in the intermediate regime $p^{d/2}\ll n\ll p^{d-1}$, where the statistical minimax rate and the best known efficient estimators do not match. More broadly, our results point to a new class of open problems in which the computational difficulty is driven not only by the statistical model, but also by the intrinsic hardness of the loss used to evaluate performance.
Acknowledgments
The authors thank Wei Biao Wu and Cheng Mao for helpful discussions.
Appendix A Additional Results for the Third Moment
In this section, we discuss more theoretical results for the third moments.
A.1 Existence
A fundamental question closely related to moment tensors concerns existence: given finitely many moments of an (unknown) probability distribution, does there exist a distribution attaining these moments, and if so, can the distribution be constructed? This question is known as the truncated moment problem. It has been extensively studied and has numerous applications, including extreme problems, optimization, and limit theorems in probability theory. General existence results can be found, for example, in Chapter 17 of [49]. In this section, we provide an explicit construction of a random vector whose first three moments match prescribed values.
Lemma 2.
For any real value $w$, there exists:
  • a random variable $X$ such that
    $\mathbb{E}X=0,\mathbb{E}X^{2}=1,\mathbb{E}X^{3}=w.$
  • two random variables $X,Y$ such that
    $\mathbb{E}X=\mathbb{E}Y=0,\quad\mathbb{E}[X,Y]^{\top}[X,Y]=\mathbf{I}_{2\times 2},\quad\mathbb{E}X^{3}=\mathbb{E}Y^{3}=\mathbb{E}X^{2}Y=0,\quad\mathbb{E}XY^{2}=w.$
Proof.
The first part of this lemma can be achieved by assuming $X$ is Bernoulli distribution. Specifically, let $\mathbb{P}(X=-ga)=1-g,\mathbb{P}(X=(1-g)a)=g$ with $g=1/2+b/2$, $a>0$ and 1<b<1-1<b<1. Obviously, $\mathbb{E}(X)=0$. As $\mathbb{E}(X^{2})=1$ and $\mathbb{E}(X^{3})=w$, we have
$\left(\frac{1}{4}-\frac{1}{4}b^{2}\right)a^{2}$ $=1,$
$-ab$ $=w.$
It follows that $b=\text{sgn}(w)(4/w^{2}+1)^{-1/2}$. Then, for any given $w$, there exists such a $b$ ensuring that the Bernoulli distribution $X$ is well-defined.
For the second part of the lemma, consider random variables $X=0,-a_{1},-a_{2},a_{3}$, $Y=0,-a_{4},a_{4}$, where all of $a_{1},a_{2},a_{3}$ are either positive or negative numbers. The joint distribution $(X,Y)$ satisfies the following distribution table:
$X$ $Y$ 0 $a_{4}$ $a_{4}$
0
$a_{1}$ $g_{1}$ $g_{1}$ $b_{2}$
$a_{2}$ $g_{2}$ $g_{2}$ $b_{2}$
$a_{3}$ $g_{3}$ $g_{3}$ $b_{3}$
$b_{1}$ $b_{1}$
Note that the final column and row in the table denote the marginal probabilities for the random variables $X$ and $Y$, respectively. To maintain degrees of freedom, some values in the distribution table have been omitted. Then, obviously $\mathbb{E}(Y)=0,\mathbb{E}(Y^{3})=0,\mathbb{E}(XY)=0,\mathbb{E}(X^{2}Y)=0$. The other conditions, $\mathbb{E}(X)=0,\mathbb{E}(X^{2})=1,\mathbb{E}(X^{3})=0,\mathbb{E}(Y^{2})=1,\mathbb{E}(XY^{2})=w$, are given by the following equations:
$-(a_{1}+a_{2})b_{2}+a_{3}b_{3}$ $=0,$ (11)
$-(a_{1}^{3}+a_{2}^{3})b_{2}+a_{3}^{3}b_{3}$ $=0,$ (12)
$(a_{1}^{2}+a_{2}^{2})b_{2}+a_{3}^{2}b_{3}$ $=1,$ (13)
$2a_{4}^{2}b_{1}$ $=1,$ (14)
$2a_{4}^{2}(a_{3}g_{3}-a_{1}g_{1}-a_{2}g_{2})$ $=w.$ (15)
Equations (11)-(13) give us
$a_{3}^{2}$ $=a_{1}^{2}+a_{2}^{2}-a_{1}a_{2},$ (16)
$b_{2}$ $=\frac{1}{(a_{1}+a_{3})(a_{2}+a_{3})},$ (17)
$b_{3}$ $=\frac{a_{1}+a_{2}}{a_{3}(a_{1}+a_{3})(a_{2}+a_{3})}.$ (18)
From (16), the values $a_{1},a_{2},a_{3}$ increase proportionally. As $b_{2},b_{3}$ are marginal probabilities with $2b_{2}+b_{3}\leq 1$, we need $|a_{1}|,|a_{2}|,|a_{3}|$ to greater than some constant (e.g. $|a_{1}|,|a_{2}|,|a_{3}|\geq 1$).
Substituting (14) into (15) and set $g_{1}=g_{2}=0$, we have
$w=a_{3}g_{3}/b_{1}.$
When $|w|\geq 1$, as $b_{3}\geq 2g_{3}$, we have $|a_{3}|\geq 2b_{1}|w|/b_{3}$ with $2b_{1}$ close to $b_{3}$. There exists such a $|a_{3}|$ ensuring that the joint distribution of $(X,Y)$ is well-defined. If $|w|<1$, then we can always set a sufficiently small $g_{3}$ to satisfy $a_{3}=wb_{1}/g_{3}$, which also ensures that the joint distribution of $(X,Y)$ is well-defined. Therefore, the second part can be constructed using the above joint discrete distribution $(X,Y)$. ∎
Theorem 10.
Suppose $\mathbf{M}\in\mathbb{R}^{p\times p}$ is a symmetric matrix and ${\mathbfcal{T}}\in\mathbb{R}^{p\times p\times p}$ is a symmetric tensor. If and only if $\mathbf{M}$ is positive semidefinite with eigenvectors associated with nonzero eigenvalues as $u_{1},\ldots,u_{r}]=\mathbf{U}$, and ${\mathbfcal{T}}\times_{1}\mathbf{U}_{\perp}=0$, there exists a random vector $X$ such that $\mathbb{E}(X)=0,\mathbb{E}(XX^{\top})=\mathbf{M},\mathbb{E}(X\circ X\circ X)={\mathbfcal{T}}$.
Proof.
Sufficiency.
We begin with the first part of the statement: for given $\mathbf{M}$ is positive semidefinite with eigenvectors associated with nonzero eigenvalues as $u_{1},\ldots,u_{m}]=\mathbf{U}$, and $\times_{1}\mathbf{U}_{\perp}=0$, there exists a random vector $X$ such that $\mathbb{E}(X)=0,\mathbb{E}(XX^{\top})=\mathbf{M},\mathbb{E}(X\circ X\circ X)={\mathbfcal{T}}$.
First, we consider a special case when $\mathbf{M}=\mathbf{I}$. Notice the following facts:
  1. 1.
    Let $X=(X_{1},\ldots,X_{p})\sim N(0,\sigma^{2}I)$ and $Y=X^{\top}\mathbf{S}X-\sigma^{2}\text{tr}(\mathbf{S})$ for some symmetric $\mathbf{S}$. Then we have $\mathbb{E}Y=\mathbb{E}X_{i}=0,\mathbb{E}X_{i}Y=\mathbb{E}X_{i}X_{j}=0,\mathbb{E}X_{i}^{2}=\sigma^{2},\mathbb{E}Y^{2}=2\sigma^{4}\text{tr}(\mathbf{S}^{2}),\mathbb{E}YX_{i}X_{j}=2\sigma^{4}\mathbf{S}_{i,j},\mathbb{E}Y^{2}X_{i}=\mathbb{E}X_{i}^{3}=0,\mathbb{E}Y^{3}=8\sigma^{6}\text{tr}(\mathbf{S}^{3})$.
  2. 2.
    For random vectors $X\perp Y$ with $\mathbb{E}X=\mathbb{E}Y=0$, we have $\mathbb{E}(X+Y)^{\otimes 3}=\mathbb{E}X^{\otimes 3}+\mathbb{E}Y^{\otimes 3}$ and $\mathbb{E}(X+Y)^{\otimes 2}=\mathbb{E}X^{\otimes 2}+\mathbb{E}Y^{\otimes 2}$.
Now, for a given tensor $\mathbfcal{T}$}, we consider the following construction procedure. For the first slice ${\mathbfcal{T}}_{1,i,j}$:
  1. 1.
    We let matrix $\mathbf{S}_{1}\in\mathbb{R}^{(p-1)\times(p-1)}$ be $(\mathbf{S}_{1})_{i,j}=({\mathbfcal{T}})_{1,1+i,1+j}$. Let $X_{1,1}=(Y,X)$ be constructed in the fact 1 with $\mathbf{S}=\mathbf{S}_{1}/2$ and $\sigma=(2\text{tr}(\mathbf{S}^{2}))^{-1/2}$.
  2. 2.
    For $k=2,\dots,p$, let $X_{1,k}$ be the vector with first entry as $(p-1)^{-1/6}X$, $k$th entry as $(p-1)^{1/3}Y$, and 0 otherwise. $X$ and $Y$ are constructed as in the part 2 of Lemma 2 such that $\mathbb{E}[X,Y]^{\top}[X,Y]=\beta\mathbf{I}_{2\times 2}$, $\mathbb{E}XY^{2}=0$, and $\mathbb{E}X^{2}Y=\sigma^{4}({\mathbfcal{T}})_{1,1,k}$ for $\beta=\sigma^{2}(p-1)^{-2/3}$.
  3. 3.
    Let $X_{1,p+1}=(X,Y)$ where $X$ is constructed as in the part 1 of Lemma 2 such that $\mathbb{E}X^{2}=\sigma^{2},\mathbb{E}X^{3}=\sigma^{4}({\mathbfcal{T}})_{1,1,1}-8\sigma^{6}\text{tr}(\mathbf{S}^{3})$, and $Y\sim N(0,\sigma^{2}\mathbf{I}_{(p-1)\times(p-1)}),Y\perp X$.
  4. 4.
    Let $\tilde{X}_{1}=\sum_{k\in[p+1]}X_{1,k}$. By the fact 2 above, we have $(\mathbb{E}\tilde{X}_{1}^{\otimes 3})_{1,i,j}=\sigma^{4}({\mathbfcal{T}})_{1,i,j}$ and $\mathbb{E}\tilde{X}_{1}^{\otimes 2}=(\sigma^{2}+\beta(p-1)^{2/3}+\sigma^{2})\mathbf{I}_{p\times p}=3\sigma^{2}\mathbf{I}_{p\times p}$.
  5. 5.
    Let $Y$ be independent of $\tilde{X}_{1}$ and constructed as in the part 1 of Lemma 2 such that $\mathbb{E}Y^{2}=1/(3p\sigma^{2})$ and $\mathbb{E}Y^{3}=1/\sigma^{4}$. Now let $X_{1}=\tilde{X}_{1}Y$. We have $(\mathbb{E}X_{1}^{\otimes 3})_{1,i,j}=({\mathbfcal{T}})_{1,i,j}$, $(\mathbb{E}X_{1}^{\otimes 3})_{k,i,j}=0$ for $i,j,k>1$, and $\mathbb{E}X_{1}^{\otimes 2}=p^{-1}\mathbf{I}$.
For the $h$th slice ${\mathbfcal{T}}_{h,i,j}$:
  1. 1.
    We let matrix $\mathbf{S}_{h}\in\mathbb{R}^{(p-h)\times(p-h)}$ be $(\mathbf{S}_{h})_{i,j}=({\mathbfcal{T}})_{h,h+i,h+j}$. Let $X_{h,h}=(Y,X)$ be constructed in the fact 1 with $\mathbf{S}=\mathbf{S}_{h}/2$ and $\sigma=(2\text{tr}(\mathbf{S}^{2}))^{-1/2}$.
  2. 2.
    For $k=h+1,\dots,p$, let $X_{h,k}$ be the vector with first entry as $(p-h)^{-1/6}X$, $k$th entry as $(p-h)^{1/3}Y$, and 0 otherwise. $X$ and $Y$ are constructed as in the part 2 of Lemma 2 such that $\mathbb{E}[X,Y]^{\top}[X,Y]=\beta\mathbf{I}_{2\times 2}$, $\mathbb{E}XY^{2}=0$, and $\mathbb{E}X^{2}Y=\sigma^{4}({\mathbfcal{T}})_{h,h,k}$ for $\beta=\sigma^{2}(p-h)^{-2/3}$.
  3. 3.
    Let $X_{h,p+1}=(X,Y)$ where $X$ is constructed as in the part 1 of Lemma 2 such that $\mathbb{E}X^{2}=\sigma^{2},\mathbb{E}X^{3}=\sigma^{4}({\mathbfcal{T}})_{h,h,h}-8\sigma^{6}\text{tr}(\mathbf{S}^{3})$, and $Y\sim N(0,\sigma^{2}\mathbf{I}_{(p-h)\times(p-h)}),Y\perp X$.
  4. 4.
    Let $\tilde{X}_{h}=\sum_{k=h}^{p+1}X_{h,k}$. By the fact 2 above, we have $(\mathbb{E}\tilde{X}_{h}^{\otimes 3})_{1,i,j}=\sigma^{4}({\mathbfcal{T}})_{h,h-1+i,h-1+j}$ and $\mathbb{E}\tilde{X}_{h}^{\otimes 2}=(\sigma^{2}+\beta(p-h)^{2/3}+\sigma^{2})\mathbf{I}_{(p+1-h)\times(p+1-h)}=3\sigma^{2}\mathbf{I}_{(p+1-h)\times(p+1-h)}$.
  5. 5.
    Let $Y$ be independent of $\tilde{X}_{h}$ and constructed as in the part 1 of Lemma 2 such that $\mathbb{E}Y^{2}=1/(3p\sigma^{2})$ ($1/(p\sigma^{2})$ if $h=p$) and $\mathbb{E}Y^{3}=1/\sigma^{4}$. Now let $\bar{X}_{h}=\tilde{X}_{h}Y$. We have $(\mathbb{E}\bar{X}_{h}^{\otimes 3})_{1,i,j}=({\mathbfcal{T}})_{h,h-1+i,h-1+j}$, $(\mathbb{E}\bar{X}_{h}^{\otimes 3})_{k,i,j}=0$ for $i,j,k>1$, and $\mathbb{E}\bar{X}_{h}^{\otimes 2}=p^{-1}\mathbf{I}_{(p+1-h)\times(p+1-h)}$.
  6. 6.
    Let $X_{h}=(0,\ldots,0,\bar{X}_{h})\in\mathbb{R}^{p}$.
We finally let $\tilde{X}=\sum_{h\in[p]}X_{h}$. We have $\mathbb{E}\tilde{X}^{\otimes 3}={\mathbfcal{T}}$ and $\mathbb{E}\tilde{X}^{\otimes 2}$ is a diagonal matrix with $k$th diagonal entry $(\mathbb{E}\tilde{X}^{\otimes 2})_{k,k}=k/p$. Let $Y\sim N(0,\mathbf{D})$ with diagonal matrix $\mathbf{D}\in\mathbb{R}^{p}$ given by $(\mathbf{D})_{kk}=(p-k)/p$. Let $X=\tilde{X}+Y$. Then we have $\mathbb{E}X^{\otimes 3}={\mathbfcal{T}}$ and $\mathbb{E}X^{\otimes 2}=\mathbf{I}$.
Now, for geneal $\mathbf{M}$, and an order-3 symmetric tensor $\mathbfcal{T}$} with full Tucker rank, if random vector $X$ satisfy $\mathbb{E}XX^{\top}=\mathbf{M}$ and $\mathbb{E}X\circ X\circ X={\mathbfcal{T}}$, then we let $Y=\mathbf{A}X$ where $\mathbf{A}$ satisfies $\mathbf{A}=(\mathbf{U}\mathbf{\Sigma}^{1/2})^{\dagger}=\mathbf{\Sigma}^{-1/2}\mathbf{U}^{\top}$ and $\mathbf{M}=\mathbf{U}\mathbf{\Sigma}\mathbf{U}^{\top}$ is the SVD. Then, it follows $\mathbb{E}YY^{\top}=\mathbf{I}$ and $\mathbb{E}Y\circ Y\circ Y={\mathbfcal{T}}\times_{1}\mathbf{A}\times_{2}\mathbf{A}\times_{3}\mathbf{A}$. Thus, we can apply the above argument to construct $Y$ and let $X=\mathbf{A}^{\dagger}Y$. So we have $\mathbb{E}XX^{\top}=\mathbf{A}^{\dagger}(\mathbf{A}^{\dagger})^{\top}=\mathbf{M}$ and $\mathbb{E}X\circ X\circ X={\mathbfcal{T}}\times_{1}\mathbf{U}\mathbf{U}^{\top}\times_{2}\mathbf{U}\mathbf{U}^{\top}\times_{3}\mathbf{U}\mathbf{U}^{\top}={\mathbfcal{T}}$.
Necessity.
Now prove the second part, i.e., for any random vector $X$ such that $\mathbb{E}(X)=0,\mathbb{E}(XX^{\top})=\mathbf{M},\mathbb{E}(X\circ X\circ X)={\mathbfcal{T}}$, we have $\mathbf{M}$ is positive semidefinite with eigenvectors associated with nonzero eigenvalues as $u_{1},\ldots,u_{m}]=\mathbf{U}$, and ${\mathbfcal{T}}\times_{1}\mathbf{U}_{\perp}=0$. The argument about the $\mathbf{M}$ is trivially true by the property of the covariance matrix. When $\operatorname{rank}(\mathbf{M})=p$, $\mathbf{U}_{\perp}$ is null so the argument about the $\mathbfcal{T}$} is trivial. When $r:=\operatorname{rank}(M_{2})<p$, we let $Y=\mathbf{A}X\in\mathbb{R}^{r}$ where $\mathbf{A}$ satisfies $\mathbf{A}=(\mathbf{U}\mathbf{\Sigma}^{1/2})^{\dagger}=\mathbf{\Sigma}^{-1/2}\mathbf{U}^{\top}$ and $\mathbf{M}=\mathbf{U}\mathbf{\Sigma}\mathbf{U}^{\top}$ is the SVD. Then, we have ${\mathbfcal{T}}=\mathbb{E}X\circ X\circ X={\mathbfcal{T}}\times_{1}\mathbf{U}\mathbf{U}^{\top}\times_{2}\mathbf{U}\mathbf{U}^{\top}\times_{3}\mathbf{U}\mathbf{U}^{\top}$, which implies the argument about $\mathbfcal{T}$} is true. ∎
A.2 A lower bound
In this section, we derive an additional lower bound for the estimation of third-order moments and cumulants, under a more restrictive and regular distributional class than that considered in Section 3.1.
Theorem 11.
Assume $p>C_{1}$ for some absolute constant $C_{1}$. Let $\mathcal{F}=\{X\in\mathbb{R}^{p}:\mathbb{E}X=0,\mathbb{E}XX^{\top}=2\mathbf{I}\}$. Then for any estimator $\hat{\mathbfcal{T}}$ from $n$ i.i.d. samples, we have
$\max_{X\in\mathcal{F}}\mathbb{E}\left\|\hat{\mathbfcal{T}}-\mathbb{E}X^{\otimes 3}\right\|\gtrsim c\sqrt{\frac{p}{n}}\land 1.$
Remark 6.
Note that the random vectors in $\mathcal{F}$ in Theorem 11 is mean 0, which implies the stated rate $\sqrt{\frac{p}{n}}\land 1$ is the minimax lower bound for the estimation of order-3 moment, central moment, and cumulant.
Proof.
We define the following Gaussian mixture model
$X\sim\operatorname{GMM}\left(\mu,(\mathbf{\Sigma}_{i,j}^{(l)})_{i,j,l\in[p]}\right)=\frac{1}{2p}\sum_{l\in[p]}N(\mu_{l},\mathbf{I}+\mathbf{\Sigma}^{(l)})+N(-\mu_{l},\mathbf{I}-\mathbf{\Sigma}^{(l)}).$
where $\mu_{l}\in\mathbb{R}^{p}$ with $(\mu_{l})_{i}=\mu$ if $i=l$ and 0 otherwise.
By Theorem 4.1 in [46], we have $\mathbb{E}X=0$, $\mathbb{E}XX^{\top}=\mathbf{I}+\frac{1}{p}\sum_{l\in[p]}\mu_{l}\mu_{l}^{\top}=(1+\frac{\mu^{2}}{p})\mathbf{I}$, and
$(\mathbb{E}X\circ X\circ X)_{i,j,k}=\frac{1}{p}\sum_{l\in[p]}(\mu_{l})_{i}\mathbf{\Sigma}^{(l)}_{j,k}+(\mu_{l})_{j}\mathbf{\Sigma}^{(l)}_{i,k}+(\mu_{l})_{k}\mathbf{\Sigma}^{(l)}_{i,j}=\frac{\mu}{p}\left(\mathbf{\Sigma}^{(i)}_{j,k}+\mathbf{\Sigma}^{(j)}_{i,k}+\mathbf{\Sigma}^{(k)}_{i,j}\right)$. (19)
Now, for all $i\in[p]$, we let $\mathbf{\Sigma}^{(i)}_{j,k}\sim N(0,\varepsilon^{2})$ for $j,k\in\{1,2\}$ and otherwise 0. Generate $\mathcal{N}=\{(\mathbf{\Sigma}^{(i)}_{j,k})_{h};h\in[N]\}$ and denote $\mathbf{\Sigma}_{h}$ as the tensor with entry
$(\mathbf{\Sigma}_{h})_{i,j,k}=\left((\mathbf{\Sigma}^{(i)}_{j,k})_{h}+(\mathbf{\Sigma}^{(j)}_{i,k})_{h}+(\mathbf{\Sigma}^{(k)}_{i,j})_{h}\right).$
Note that the entries $(i,j,k)$ of $\mathbf{\Sigma}_{h}-\mathbf{\Sigma}_{q}$ with $i\geq 2,j,k\in\{1,2\}$ are i.i.d. from $N(0,2\varepsilon^{2})$. Thus, we have
$\mathbb{P}\left(\|\mathbf{\Sigma}_{h}-\mathbf{\Sigma}_{q}\|<c\varepsilon\sqrt{p}\right)$
$\leq$ $\mathbb{P}\left(z<c\varepsilon\sqrt{p-2}\right)$
$<$ $2\exp\left(-C(p-2)\right)$
$<$ $2\exp\left(-Cp\right),$
where $z\sim N(0,2\varepsilon^{2}\mathbf{I})\in\mathbb{R}^{p-2}$.
Further note that $\|(\mathbf{\Sigma}^{(i)})_{h}\|\leq 2|(\mathbf{\Sigma}^{(i)}_{1,2})_{h}|\sim N(0,4\varepsilon^{2})$ and hence $\mathbb{P}\left(\sum_{i\in[p]}\|(\mathbf{\Sigma}^{(i)})_{h}\|/p>c\varepsilon\right)\leq\exp(-Cp)$. Thus, by the union bound argument, we can generate $\mathcal{N}$ with $N=e^{Cp}$, such that for all $h,q\in[N]$, we have
$\|\mathbf{\Sigma}_{h}-\mathbf{\Sigma}_{q}\|>c\varepsilon\sqrt{p},$ (20)
$\operatorname{rank}((\mathbf{\Sigma}^{(l)})_{h})=2,$
and
$\sum_{l\in[p]}\|(\mathbf{\Sigma}^{(l)})_{h}\|/p<C\varepsilon.$
Hence, we have $\sum_{l\in[p]}\|(\mathbf{\Sigma}^{(l)})_{h}\|_{F}/p<C\varepsilon$. Thus, for $(\mathbf{\Sigma}_{i,j}^{(l)})_{i,j,l\in[p]},(\tilde{\mathbf{\Sigma}}_{i,j}^{(l)})_{i,j,l\in[p]}\in\mathcal{N}$, by Lemma 8 and 9, we have
$D\left(\operatorname{GMM}\left(\mu,(\mathbf{\Sigma}_{i,j}^{(l)})_{i,j,l\in[p]}\right)\|\operatorname{GMM}\left(\mu,(\tilde{\mathbf{\Sigma}}_{i,j}^{(l)})_{i,j,l\in[p]}\right)\right)$ $\leq$ $C\varepsilon^{2},$ (21)
for $\varepsilon<c$.
Note that if $X\sim\operatorname{GMM}\left(\mu,(\mathbf{\Sigma}_{i,j}^{(l)})_{i,j,l\in[p]}\right)$ and $Y\sim\operatorname{GMM}\left(\mu,(\tilde{\mathbf{\Sigma}}_{i,j}^{(l)})_{i,j,l\in[p]}\right)$, then by (20), we have
$\|\mathbb{E}X\circ X\circ X-\mathbb{E}Y\circ Y\circ Y\|\geq c\mu\varepsilon p^{-1/2}.$ (22)
Finally, by (21), (22), and Lemma 7, we have
$\max_{j}E_{\operatorname{GMM}_{j}}\left\|\hat{\theta},\theta\left(\operatorname{GMM}_{j}\right)\right\|\gtrsim\mu\varepsilon p^{-1/2}\left(1-\frac{C_{1}n\varepsilon^{2}+\log 2}{C_{2}p}\right),$
where $\theta\left(\operatorname{GMM}_{j}\right)=\mathbb{E}X\circ X\circ X$ and $X\sim\operatorname{GMM}_{j}$.
When $n>p$, we pick $\mu=\sqrt{p}$ and $\varepsilon=c\sqrt{p/n}<1$. Then we have $\mathbb{E}XX^{\top}=2\mathbf{I}$. Moreover, the lower bound follows:
$\max_{j}E_{\operatorname{GMM}_{j}}\left\|\hat{\theta}-\theta\left(\operatorname{GMM}_{j}\right)\right\|\gtrsim\sqrt{\frac{p}{n}}.$
When $n\leq p$, we let $\mu=\sqrt{p}$ and $\varepsilon=c$. Then we have $\mathbb{E}XX^{\top}=2\mathbf{I}$. Moreover, the lower bound follows:
$\max_{j}E_{\operatorname{GMM}_{j}}\left\|\hat{\theta}-\theta\left(\operatorname{GMM}_{j}\right)\right\|\gtrsim 1.$
Appendix B Proofs
B.1 Proof of Theorem 2
For moments:
Proof.
Consider random vectors $X\sim N(\mu,\mathbf{I})\in\mathbb{R}^{p}$. Then by Theorem 3.1 in [46], We have $\mathbb{E}X^{\otimes d}=\sum_{k=0}^{\lfloor d/2\rfloor}K_{d,k}\operatorname{sym}\left(\mu^{\otimes d-2k}\otimes\mathbf{I}^{\otimes k}\right)$ with $K_{d,k}=\binom{d}{2k}\frac{(2k)!}{k!2^{k}}$ and $\operatorname{sym}$ defined as in [46]. We generate $N=\exp(Cp)$ of vectors $\mu_{1}\ldots\mu_{N}\in\mathbb{R}^{p}$ such that $\|\mu_{i}\|=\mu$, $\|\mu_{i}-\mu_{j}\|>0.1\mu$ and $\langle\mu_{i},\mu_{j}\rangle>0$ for any $i,j\in[N]$. By Lemma 4, we have
$\|\mu_{i}^{\otimes d-2k}\otimes\mathbf{I}^{\otimes k}-\mu_{j}^{\otimes d-2k}\otimes\mathbf{I}^{\otimes k}\|\geq\|\mu_{i}^{\otimes d-2k}-\mu_{j}^{\otimes d-2k}\|\geq\|\mu_{i}-\mu_{j}\|/2,$
which further implies
$\|\mathbb{E}X^{\otimes d}-\mathbb{E}Y^{\otimes d}\|=\sum_{k=0}^{\lfloor d/2\rfloor}K_{d,k}\left(\operatorname{sym}\left(\mu_{i}^{\otimes d-2k}\otimes\mathbf{I}^{\otimes k}\right)-\operatorname{sym}\left(\mu_{j}^{\otimes d-2k}\otimes\mathbf{I}^{\otimes k}\right)\right)\geq C_{d}\|\mu_{i}-\mu_{j}\|$
for $X\sim N(\mu_{i},\mathbf{I})$ and $Y\sim N(\mu_{j},\mathbf{I})$. We also have
$D\left(X\|Y\right)=\|\mu_{i}-\mu_{j}\|^{2}.$
Thus, Lemma 7 implies
$\max_{X\sim N(\mu_{i},I),i\in[N]}\mathbb{E}_{X}\left\|\hat{\mathbfcal{T}}-\mathbb{E}X^{\otimes d}\right\|\geq C_{d}\mu\left(1-\frac{c_{d}n\mu^{2}+\log 2}{p}\right).$
Let $\mu=\sqrt{(n\land p)/n}$. Then we have $\|\mu_{i}\|=\mu<1$ so that $\|v^{\top}X\|_{\Phi_{2}}<C$ for some absolute constant $C$, Moreover, we have
$\max_{X\sim N(\mu_{i},I),i\in[N]}\mathbb{E}_{X}\left\|\hat{\mathbfcal{T}}-\mathbb{E}X^{\otimes d}\right\|\geq C_{d}\sqrt{p/n}\land 1.$
For cumulants:
Proof.
We assume the following claim and prove it later:
Claim: There exist random variables $X_{u,\theta}$ for $\theta\in(0,\theta_{0}]$ and $u\in\mathbb{S}^{p-1}$ for some constant $\theta_{0}$ which does not depend on $p$ and $u$, such that
$\sup_{u\in\mathbb{S}^{p-1},\theta\in(0,\theta_{0}]}\|X_{u,\theta}\|_{\Phi_{2}}\leq 2,$ (23)
$\|{\mathbfcal{K}}_{d}(X_{u,\theta})-{\mathbfcal{K}}_{d}(X_{v,\theta})\|\geq C_{1}\theta\|u^{\otimes d}-v^{\otimes d}\|-C_{2}\theta^{2},$ (24)
and
$\sup_{u\in\mathbb{S}^{p-1},\theta\in(0,\theta_{0}]}D(X_{u,\theta}||X_{v,\theta})\leq C_{3}\theta^{2}\|u-v\|^{2}+C_{4}\theta^{3},$ (25)
where $C_{1},C_{2},C_{3}$, and $C_{4}$ are constants that only depend on $d$.
Now we prove the theorem. For some $r>\exp(cp)$, there exist $\mathcal{U}=\{u_{1},\ldots,u_{r}\}$ such that $\|u_{i}-u_{j}\|>1/2$. We construct the following family of random vectors for $\theta\in(0,\theta_{0}]$:
$\mathcal{F}_{\theta}=\{X_{u,\theta}:u\in\mathcal{U}\}.$
Now, combining Lemma 4, Lemma 7, (23), (24), and (25), it follows that for any $\theta\in(0,\theta_{0}]$,
$\inf_{\hat{\mathbfcal{K}}}\max_{X\in\mathcal{F}}\mathbb{E}(\hat{\mathbfcal{K}}-{\mathbfcal{K}}_{d}(X))$
$\geq$ $\inf_{\hat{\mathbfcal{K}}}\max_{X\in\mathcal{F}_{\theta}}\mathbb{E}(\hat{\mathbfcal{K}}-{\mathbfcal{K}}_{d}(X))$
$\geq$ $(C_{1}\theta\|u^{\otimes d}-v^{\otimes d}\|-C_{2}\theta^{2})\left(1-\frac{nC_{3}\theta^{2}\|u-v\|^{2}+C_{4}\theta^{3}+\log 2}{cp}\right)$
$\geq$ $\left(\frac{1}{2}C_{1}\theta\|u-v\|-C_{2}\theta^{2}\right)\left(1-\frac{nC_{3}\theta^{2}\|u-v\|^{2}+C_{4}\theta^{3}+\log 2}{cp}\right)$
$\geq$ $\left(2^{-2}C_{1}\theta-C_{2}\theta^{2}\right)\left(1-\frac{n2^{-2}C_{3}\theta^{2}+C_{4}\theta^{3}+\log 2}{cp}\right).$
We pick $\theta=c_{0}(1\land\sqrt{p/n})$ for some small enough $c_{0}$. The desired result follows.
Proof of the Claim: Now we prove the claim by constructing $X_{u,\theta}$ for $\theta\in(0,\theta_{0}]$ and $u\in\mathbb{S}^{p-1}$ and prove (23), (24), and (25).
Let $H_{d}$ be the probabilists’ Hermite polynomial of degree $d$. For some $M>0$, define
$h_{M}(t):=H_{d}(t)I_{\left\{\left|H_{d}(t)\right|\leq M\right\}}+M\operatorname{sign}(H_{d}(t))I_{\left\{\left|H_{d}(t)\right|>M\right\}}.$
Let $Z\sim P_{0}={N}\left(0,I_{p}\right)$ and, for each $u\in\mathbb{S}^{p-1}$ and $\theta\in\mathbb{R}$, define the tilted distribution $P_{u,\theta}$ by
$\frac{dP_{u,\theta}}{dP_{0}}(x)=\frac{\exp\left(\theta h_{M}\left(u^{\top}x\right)\right)}{\mathbb{E}_{P_{0}}\exp\left(\theta h_{M}\left(u^{\top}Z\right)\right)}.$
We let $X_{u,\theta}\sim P_{u,\theta}$.
Proof of (23). Note that for any $s\in\mathbb{R}$ and $v\in\mathbb{S}^{p-1}$, we have
$\mathbb{E}_{P_{u,\theta}}e^{sv^{\top}X}=\frac{\mathbb{E}_{P_{0}}\exp\left(\theta h_{M}\left(u^{\top}Z\right)+sv^{\top}Z\right)}{\mathbb{E}_{P_{0}}\exp\left(\theta h_{M}\left(u^{\top}Z\right)\right)}\leq\frac{e^{\theta M}\mathbb{E}_{P_{0}}e^{sv^{\top}Z}}{e^{-\theta M}}=e^{2\theta M}e^{s^{2}/2}.$
Pick $\theta_{0,M}$ so that $e^{2\left|\theta_{0,M}\right|M}\leq e^{1/2}$. Then we have $\|v^{\top}X_{u,\theta}\|_{\psi_{2}}\leq 2$, which proves (23) as long as $M$ is a constant only depending on $d$.
Proof of (24). We decompose random vector $X_{u,\theta}$ as $X_{u,\theta}=Su+Z$ where $S=u^{\top}X_{u,\theta}\in\mathbb{R}$ and $Z=(I-uu^{\top})X_{u,\theta}\perp u$.
By calculating the joint density (with respect to Lebesgue measure) of $(S,Z)$, we can see that $Z\perp S$, $Z\sim N(0,(I-uu^{\top}))$, and
$f_{S}(s)=e^{\theta h_{M}(s)}\phi(s)/\mathbb{E}_{U\sim N(0,1)}e^{\theta h_{M}(U)}.$
For $t\in\mathbb{R}^{p}$, we decompose $t=t_{u}u+t_{\perp}$. Then the cumulant generating function of $X_{u,\theta}$ is
$\kappa_{X}(t)=\log\mathbb{E}\exp(t^{\top}X_{u,\theta})=\log\mathbb{E}\exp(t_{u}S)+\log\mathbb{E}\exp(t_{\perp}^{\top}Z).$
Note that by $Z\sim{N}(0,(I-uu^{\top}))$, we have $\mathbb{E}e^{t_{\perp}^{\top}Z}=e^{\frac{1}{2}t_{\perp}^{\top}(I-uu^{\top})t_{\perp}}=e^{\frac{1}{2}\left\|t_{\perp}\right\|^{2}}$. Thus,
$\kappa_{X}(t)=\kappa_{S}(t_{u})+{\frac{1}{2}\left\|t_{\perp}\right\|^{2}}.$
For $d\geq 3$, by the chain rule of derivatives, we have
${\mathbfcal{K}}_{d}(X_{u,\theta})=\nabla^{d}\left[\kappa_{S}(\langle t,u\rangle)\right]|_{t=0}=\kappa_{S}^{(d)}(0)u^{\otimes d}.$
Recall that $f_{S}(s)=e^{\theta h_{M}(s)}\phi(s)/\mathbb{E}_{U\sim N(0,1)}e^{\theta h_{M}(U)}$, which implies
$\kappa_{S}(t)=\log\mathbb{E}_{U\sim N(0,1)}\exp(\theta h_{M}(U)+tU)-\log\mathbb{E}_{U\sim N(0,1)}\exp(\theta h_{M}(U)).$
Thus,
$\frac{\partial\kappa_{S}(t)}{\partial\theta}=\frac{\mathbb{E}h_{M}(U)\exp(\theta h_{M}(U)+tU)}{\mathbb{E}\exp(\theta h_{M}(U)+tU)}-\frac{\mathbb{E}h_{M}(U)\exp(\theta h_{M}(U))}{\mathbb{E}\exp(\theta h_{M}(U))}.$
At $\theta=0$, we have
$\kappa_{S}(t)|_{\theta=0}=\log\mathbb{E}\exp(tU),$
and
$\frac{\partial\kappa_{S}(t)}{\partial\theta}|_{\theta=0}=\frac{\mathbb{E}h_{M}(U)\exp(tU)}{\mathbb{E}\exp(tU)}-\mathbb{E}h_{M}(U).$
Additionally, we have
$\frac{d^{d}}{dt^{d}}\log\mathbb{E}\exp(tU)|_{t=0}=0,\quad\text{for $d\geq 3$},$
and
$\frac{d^{d}}{dt^{d}}\frac{\mathbb{E}h_{M}(U)\exp(tU)}{\mathbb{E}\exp(tU)}|_{t=0}$
$=$ $\frac{d^{d}}{dt^{d}}({\mathbb{E}h_{M}(U)\exp(tU-t^{2}/2)})|_{t=0}$
$=$ $\mathbb{E}h_{M}(U)\frac{d^{d}}{dt^{d}}(\exp(tU-t^{2}/2))|_{t=0}$
$=$ $\mathbb{E}h_{M}(U)H_{d}(U)$
$:=$ $C_{M,d}.$
Note that $C_{M,d}=\mathbb{E}(H_{d}(U)^{2}I_{\{H_{d}(U)<M\}}+M\operatorname{sign}(H_{d}(U))H_{d}(U))\rightarrow\mathbb{E}H_{d}(U)^{2}$ as $M\rightarrow\infty$, so that for large enough $M$ which only depends on $d$, $C_{M,d}$ is greater than 0. By Taylor’s expansion, we have
${\mathbfcal{K}}_{d}(X_{u,\theta})$ $=$ $\frac{\partial^{d}}{\partial t^{d}}\kappa_{S}(t)|_{t=0}u^{\otimes d}$ (26)
$=$ $u^{\otimes d}\left(\theta\frac{\partial^{d+1}}{\partial\theta\partial t^{d}}\kappa_{S}(t)|_{t=\theta=0}+O(\theta^{2})\right)$
$=$ $u^{\otimes d}\left(\theta\frac{\partial^{d+1}}{\partial t^{d}\partial\theta}\kappa_{S}(t)|_{t=\theta=0}+O(\theta^{2})\right)$
$=$ $(C_{M,d}\theta+O(\theta^{2}))u^{\otimes d}.$
Thus, this implies (24).
Proof of (25). Denote $A_{u}(\theta)=\log\mathbb{E}_{U\sim N(0,I)}\exp(\theta h_{M}(u^{\top}U))$. We have
$D(X_{u,\theta}||X_{v,\theta})=\mathbb{E}_{X\sim P_{u,\theta}}\theta(h_{M}(u^{\top}X)-h_{M}(v^{\top}X))-A_{u}(\theta)+A_{v}(\theta).$
By Taylor expansion at $\theta=0$, we have
$A_{u}(\theta)=\theta\mathbb{E}h_{M}(u^{\top}U)+0.5\theta^{2}\text{Var}(h_{M}(u^{\top}U))+O(\theta^{3}),$
$\phi_{u,v}(\theta)$ $:=$ $\mathbb{E}_{X\sim P_{u,\theta}}h_{M}(v^{\top}X)$
$=$ $\frac{\mathbb{E}h_{M}(v^{\top}U)e^{\theta h_{M}(u^{\top}U)}}{\mathbb{E}e^{\theta h_{M}(u^{\top}U)}}$
$=$ $\mathbb{E}h_{M}(v^{\top}U)+\theta\text{Cov}\left(h_{M}(v^{\top}U),h_{M}(u^{\top}U)\right)+O(\theta^{2}).$
and
$\phi_{u,u}(\theta)=\mathbb{E}h_{M}(u^{\top}U)+\theta\text{Var}\left(h_{M}(u^{\top}U)\right)+O(\theta^{2}).$
Hence, we have
$D(X_{u,\theta}||X_{v,\theta})$ $=$ $\frac{1}{2}\theta^{2}\text{Var}\left(h_{M}(u^{\top}U)-h_{M}(v^{\top}U)\right)+O(\theta^{2}).$
Further note that there exist some $L_{d,M}$ such that $h_{M}(a)-h_{M}(b)\leq L_{d,M}|a-b|$. Thus, we have
$\text{Var}\left(h_{M}(u^{\top}U)-h_{M}(v^{\top}U)\right)\leq L_{d,M}^{2}\mathbb{E}\left(U^{\top}(u-v)\right)^{2}=L_{d,M}^{2}\|u-v\|^{2}.$
Thus, we finally have
$D(X_{u,\theta}||X_{v,\theta})\leq C_{d,M}^{2}\theta^{2}\|u-v\|^{2}+O(\theta^{3}),$
which proves (25).
B.2 Proof of Theorem 3
Proof.
We use the standard testing form of Fano’s lemma: there exist absolute constants $\alpha_{0},c_{0}>0$ such that if probability measures $Q_{0},Q_{1},\ldots,Q_{N}$ satisfy
$\frac{1}{N}\sum_{i=1}^{N}D(Q_{i}\|Q_{0})\leq\alpha_{0}\log N,$
then
$\inf_{\psi}\left\{Q_{0}(\psi=1)+\frac{1}{N}\sum_{i=1}^{N}Q_{i}(\psi=0)\right\}\geq c_{0}.$
In particular,
$\inf_{\psi}\left\{Q_{0}(\psi=1)+\max_{1\leq i\leq N}Q_{i}(\psi=0)\right\}\geq c_{0}.$
We prove the cumulant and moment cases separately.
Cumulants. Let $P_{0}$ be the law of $Z\sim N(0,I_{p})$. Since $d\geq 3$, we have ${\mathbfcal{K}}_{d}(Z)=0$, and clearly $P_{0}\in\mathcal{F}$.
Next, let $\mathcal{U}=\{u_{1},\ldots,u_{N}\}\subset\mathbb{S}^{p-1}$ be such that
$\log N\geq cp$
for some absolute constant $c>0$; for example, one may take a maximal $1/2$-packing of $\mathbb{S}^{p-1}$.
Now use the tilted family $X_{u,\theta}$ constructed in the proof of Theorem 2. By the proof of (23), for every $v\in\mathbb{S}^{p-1}$ we have $|v^{\top}X_{u,\theta}|{\Phi_{2}}\leq 2$, hence $X_{u,\theta}\in\mathcal{F}$. Moreover, by (26),
${\mathbfcal{K}}_{d}(X_{u,\theta})=\big(C_{M,d}\theta+O(\theta^{2})\big)u^{\otimes d},$
where $C_{M,d}>0$ once $M=M(d)$ is chosen sufficiently large. Since $\|u^{\otimes d}\|=1$, it follows that for all sufficiently small $\theta$,
$\|{\mathbfcal{K}}_{d}(X_{u,\theta})\|\geq c_{1}\theta$
for some constant $c_{1}>0$ depending only on $d$. Therefore, if we set
$\rho_{*}:=\frac{c_{1}}{2}\theta,$
then
$X_{u_{i},\theta}\in H_{1}^{(K)}(\rho_{*}),\qquad i=1,\ldots,N.$
It remains to bound the KL divergence to the null. By construction,
$\frac{dP_{u,\theta}}{dP_{0}}(x)=\frac{\exp\big(\theta h_{M}(u^{\top}x)\big)}{\mathbb{E}_{P_{0}}\exp\big(\theta h_{M}(u^{\top}Z)\big)}.$
Let
$A(\theta):=\log\mathbb{E}_{U\sim N(0,1)}e^{\theta h_{M}(U)}.$
Since $u^{\top}Z\sim N(0,1)$, this does not depend on $u$. Also,
$\log\frac{dP_{u,\theta}}{dP_{0}}(x)=\theta h_{M}(u^{\top}x)-A(\theta).$
Hence
$D(P_{u,\theta}\|P_{0})=\mathbb{E}_{P_{u,\theta}}\big[\theta h_{M}(u^{\top}X)-A(\theta)\big]=\theta A^{\prime}(\theta)-A(\theta).$
Because $h_{M}$ is bounded, $A$ is analytic near 0, so
$A(\theta)=A^{\prime}(0)\theta+\frac{1}{2}A^{\prime\prime}(0)\theta^{2}+O(\theta^{3}),\qquad A^{\prime}(\theta)=A^{\prime}(0)+A^{\prime\prime}(0)\theta+O(\theta^{2}).$
Therefore,
$D(P_{u,\theta}\|P_{0})=\frac{1}{2}A^{\prime\prime}(0)\theta^{2}+O(\theta^{3})\leq C_{2}\theta^{2}$
for all sufficiently small $\theta$, uniformly over $u\in\mathbb{S}^{p-1}$. By tensorization,
$D(P_{u,\theta}^{\otimes n}\|P_{0}^{\otimes n})=nD(P_{u,\theta}\|P_{0})\leq C_{2}n\theta^{2}.$
Choose
$\theta=c_{2}\Big(\sqrt{p/n}\wedge 1\Big)$
with $c_{2}>0$ sufficiently small so that $\theta\leq\theta_{0}$ and
$C_{2}n\theta^{2}\leq\alpha_{0}\log N.$
This is possible since $\log N\geq cp$. Hence
$\frac{1}{N}\sum_{i=1}^{N}D(P_{u_{i},\theta}^{\otimes n}\|P_{0}^{\otimes n})\leq\alpha_{0}\log N.$
Applying Fano’s lemma, we obtain
$\inf_{\psi}\left\{P_{0}^{\otimes n}(\psi=1)+\max_{1\leq i\leq N}P_{u_{i},\theta}^{\otimes n}(\psi=0)\right\}\geq c_{0}.$
Since $P_{0}\in H_{0}^{(K)}$ and $P_{u_{i},\theta}\in H_{1}^{(K)}(\rho_{*})$, it follows that
$\mathfrak{R}_{n}^{(K)}(\rho_{*})\geq c_{0}.$
Finally, if $\rho\leq\rho_{*}$, then $H_{1}^{(K)}(\rho_{*})\subseteq H_{1}^{(K)}(\rho)$, so the testing problem with threshold $\rho$ is harder. Therefore
$\mathfrak{R}_{n}^{(K)}(\rho)\geq\mathfrak{R}_{n}^{(K)}(\rho_{*})\geq c_{0}.$
Since $\rho_{*}\asymp\theta\asymp\sqrt{p/n}\wedge 1$, this proves the cumulant claim.
Moments (odd $d$). Assume now that $d$ is odd. Again let $P_{0}$ be the law of $Z\sim N(0,I_{p})$. Since $Z$ is symmetric and $d$ is odd,
$\mathbb{E}Z^{\otimes d}=0,$
so $P_{0}\in H_{0}^{(M)}$.
We use the same tilted family $X_{u,\theta}$. Fix $u\in\mathbb{S}^{p-1}$, and let
$S:=u^{\top}X_{u,\theta}.$
From the construction in the proof of Theorem 2, $S$ has density
$f_{S}(s)=\frac{e^{\theta h_{M}(s)}\phi(s)}{\mathbb{E}_{U\sim N(0,1)}e^{\theta h_{M}(U)}},$
where $\phi$ is the standard Gaussian density. Therefore,
$\mathbb{E}S^{d}=\frac{\mathbb{E}\big[U^{d}e^{\theta h_{M}(U)}\big]}{\mathbb{E}e^{\theta h_{M}(U)}}.$
Since $h_{M}$ is bounded, the numerator and denominator are analytic in $\theta$ near 0, and because $d$ is odd, $\mathbb{E}U^{d}=0$. Hence
$\mathbb{E}S^{d}=\theta\mathbb{E}\big[U^{d}h_{M}(U)\big]+O(\theta^{2}).$
Also, $h_{M}\to H_{d}$ as $M\to\infty$, and
$\mathbb{E}\big[U^{d}H_{d}(U)\big]=d!>0.$
Thus, by choosing $M=M(d)$ sufficiently large, we have
$B_{M,d}:=\mathbb{E}\big[U^{d}h_{M}(U)\big]>0.$
Therefore
$\mathbb{E}S^{d}=B_{M,d}\theta+O(\theta^{2}),$
and for all sufficiently small $\theta$,
$|\mathbb{E}S^{d}|\geq c_{3}\theta$
for some constant $c_{3}>0$ depending only on $d$.
Now by the definition of tensor spectral norm,
$\|\mathbb{E}X_{u,\theta}^{\otimes d}\|=\sup_{\|v_{1}\|=\cdots=\|v_{d}\|=1}\left|\left\langle\mathbb{E}X_{u,\theta}^{\otimes d},\,v_{1}\otimes\cdots\otimes v_{d}\right\rangle\right|.$
Choosing $v_{1}=\cdots=v_{d}=u$ gives
$\|\mathbb{E}X_{u,\theta}^{\otimes d}\|\geq\left|\left\langle\mathbb{E}X_{u,\theta}^{\otimes d},\,u^{\otimes d}\right\rangle\right|=|\mathbb{E}\langle u,X_{u,\theta}\rangle^{d}|=|\mathbb{E}S^{d}|\geq c_{3}\theta.$
Hence, if we define
$\rho_{*}^{\prime}:=\frac{c_{3}}{2}\theta,$
then
$X_{u_{i},\theta}\in H_{1}^{(M)}(\rho_{*}^{\prime}),\qquad i=1,\ldots,N.$
The KL bound to the null is exactly the same as above:
$D(P_{u,\theta}^{\otimes n}\|P_{0}^{\otimes n})\leq C_{2}n\theta^{2}.$
With the same choice
$\theta=c_{2}\Big(\sqrt{p/n}\wedge 1\Big),$
Fano’s lemma yields
$\mathfrak{R}_{n}^{(M)}(\rho_{*}^{\prime})\geq c_{0}.$
By monotonicity in $\rho$, for every $\rho\leq\rho_{*}^{\prime}$,
$\mathfrak{R}_{n}^{(M)}(\rho)\geq\mathfrak{R}_{n}^{(M)}(\rho_{*}^{\prime})\geq c_{0}.$
Since $\rho_{*}^{\prime}\asymp\theta\asymp\sqrt{p/n}\wedge 1$, this proves the moment claim.
Setting $\eta_{d}=\eta_{d}^{\prime}:=c_{0}$ and adjusting the constants $c_{d},c_{d}^{\prime}>0$ completes the proof. ∎
B.3 Proof of Theorem 4
We first prove the result for the raw moments.
Proof.
for the raw moments.
By Lemma 5, we have
$\left\|\frac{1}{n}\sum_{i=1}^{n}X_{i}^{\otimes d}-\mathbb{E}(X_{i}^{\otimes d})\right\|$
[…]
$=$ $\sup_{u\in\mathbb{S}^{p-1}}\left|\frac{1}{n}\sum_{i=1}^{n}(X_{i}^{\top}u)^{d}-\mathbb{E}(X_{i}^{\top}u)^{d}\right|.$
For any given $u$, we know that $X_{i}^{\top}u$ is $\beta$-subexponential random vector. By definition, we have $\mathbb{E}\exp(|u^{\top}X|^{\beta}/t^{\beta})\leq 2$ for $t\geq\|v^{\top}X\|_{\Phi_{\beta}}$, which implies $\mathbb{E}\exp(|(u^{\top}X)^{d}|^{\beta/d}/t^{\beta})\leq 2$. Thus, $(X_{i}^{\top}u)^{d}$ is $\beta/d$-subexponential random vector with $\|(X_{i}^{\top}u)^{d}\|_{\Phi_{\beta/d}}=\|X_{i}^{\top}u\|_{\Phi_{\beta}}^{d}\leq 1$, which further implies $(X_{i}^{\top}u)^{d}-\mathbb{E}(X_{i}^{\top}u)^{d}$ is a $\beta/d$-subexponential random vector. By Lemma A.1 and A.3 in [21], we have $\|(X_{i}^{\top}u)^{d}-\mathbb{E}(X_{i}^{\top}u)^{d}\|_{\Phi_{\beta/d}}\leq M_{\beta,d}\|X_{i}^{\top}u\|_{\Phi_{\beta}}^{d}\leq M_{\beta,d}$.
Hence, when $\beta\leq d$, by Lemma 6, we have
$\mathbb{P}\left(\left|\sum_{i=1}^{n}(X_{i}^{\top}u)^{d}-\mathbb{E}(X_{i}^{\top}u)^{d}\right|>t\right)\leq 2\exp\left(-C_{\beta,d}\min\left\{t^{2}/n,t^{\beta/d}\right\}\right).$
Note that the same bound hold for $\beta>d$, whereas $(X_{i}^{\top}u)^{d}-\mathbb{E}(X_{i}^{\top}u)^{d}$ is subexponential.
We next apply the $\varepsilon$-net arguments. Notice that
[…]
$\leq$ $\left\|\frac{1}{n}\sum_{i=1}^{n}X_{i}^{\otimes d}-\mathbb{E}(X_{i}^{\otimes d})\right\|\|u^{\otimes d}-v^{\otimes d}\|_{*}$
where $\|\cdot\|_{*}$ is the tensor nuclear norm, which is the dual norm of the tensor spectral norm. Additionally, for $\|u\|=\|v\|=1$, we have
$\|u^{\otimes d}-v^{\otimes d}\|_{*}$ $\leq$ $\|(u-v)\circ u\circ\ldots\circ u\|_{*}+\ldots+\|v\circ v\circ\ldots\circ(u-v)\|_{*}$
$\leq$ $d\|u-v\|$
Let $\mathcal{N}$ be an $\varepsilon$-net of $\mathbb{S}^{p-1}$ with $\varepsilon=1/(2d)$ and $|\mathcal{N}|<\exp(C_{d}p)$. Then by union bound, we have
$\mathbb{P}\left(\sup_{u\in\mathcal{N}}\frac{1}{n}\left|\sum_{i=1}^{n}(X_{i}^{\top}u)^{d}-\mathbb{E}(X_{i}^{\top}u)^{d}\right|>t\right)\leq 2\exp\left(C_{d}p-C_{\beta,d}\min\left\{t^{2}/n,t^{\beta/d}\right\}\right).$
Further note that for any $u\in\mathbb{S}^{p-1}$, there exists $v\in\mathcal{N}$ such that $\|u-v\|\leq\varepsilon=1/(2d)$. Thus, it follows that
$\left|\frac{1}{n}\sum_{i=1}^{n}(X_{i}^{\top}u)^{d}-\mathbb{E}(X_{i}^{\top}u)^{d}\right|$
[…]
which by taking the maximum over $u\in\mathbb{S}^{p-1}$ further implies
$\sup_{u\in\mathbb{S}^{p-1}}\frac{1}{n}\left|\sum_{i=1}^{n}(X_{i}^{\top}u)^{d}-\mathbb{E}(X_{i}^{\top}u)^{d}\right|\leq 2\sup_{u\in\mathcal{N}}\frac{1}{n}\left|\sum_{i=1}^{n}(X_{i}^{\top}u)^{d}-\mathbb{E}(X_{i}^{\top}u)^{d}\right|.$
Hence, we have
$\mathbb{P}\left(\sup_{u\in\mathbb{S}^{p-1}}\frac{1}{n}\left|\sum_{i=1}^{n}(X_{i}^{\top}u)^{d}-\mathbb{E}(X_{i}^{\top}u)^{d}\right|>t\right)\leq 2\exp\left(C_{d}p-C_{\beta,d}\min\left\{t^{2}n,(tn)^{\beta/d}\right\}\right).$
Take $t=C_{\beta,d}\max\{p^{d/\beta}/n,\sqrt{p/n}\}$, the upper bound for the raw moments follows. ∎
For simplicity, we prove the upper bound for the central moments when $d=3$. For general $d$, one only needs to replace (27) with the order-$d$ expansion and bound each term using the same arguments as in the following proof.
Proof.
for the central moments.
By Lemma 5, we have
$\left\|\frac{1}{n}\sum_{i=1}^{n}(X_{i}-\bar{X})^{\otimes 3}-\mathbb{E}(X_{i}-\mu)^{\otimes 3}\right\|$ (27)
$=$ $\sup_{u\in\mathbb{S}^{p-1}}\left|\frac{1}{n}\sum_{i=1}^{n}((X_{i}-\bar{X})^{\top}u)^{3}-\mathbb{E}((X_{i}-\mu)^{\top}u)^{3}\right|$
$\leq$ $\sup_{u\in\mathbb{S}^{p-1}}\left|\frac{1}{n}\sum_{i=1}^{n}((X_{i}-\mu)^{\top}u)^{3}-\mathbb{E}((X_{i}-\mu)^{\top}u)^{3}\right|$
$+3\sup_{u\in\mathbb{S}^{p-1}}((\bar{X}-\mu)^{\top}u)\frac{1}{n}\sum_{i=1}^{n}((X_{i}-\mu)^{\top}u)^{2}$
$+2\sup_{u\in\mathbb{S}^{p-1}}((\bar{X}-\mu)^{\top}u)^{3}.$
The first term can be controlled by the arguments used in the proof for the raw moments. Following the same arguments, we also have
$\mathbb{P}\left(\sup_{u\in\mathbb{S}^{p-1}}|(\bar{X}-\mu)^{\top}u|\leq C_{\beta}\max\{\sqrt{p/n},(p/n)^{1/\beta}\}\right)\geq 1-\exp(-c_{\beta}p),$
and
$\mathbb{P}\left(\sup_{u\in\mathbb{S}^{p-1}}\left|\frac{1}{n}\sum_{i=1}^{n}((X_{i}-\mu)^{\top}u)^{2}-\mathbb{E}((X_{i}-\mu)^{\top}u)^{2}\right|\leq C_{\beta}\max\{\sqrt{p/n},p^{2/\beta}/n\}\right)$
$\geq$ $1-\exp(-c_{\beta}p).$
Moverover, we have $\sup_{u\in\mathbb{S}^{p-1}}\mathbb{E}((X_{i}-\mu)^{\top}u)^{2}\leq C_{\beta}$. Thus, with probability greater than $1-\exp(-c_{\beta}p)$, we have
$\sup_{u\in\mathbb{S}^{p-1}}\frac{1}{n}\sum_{i=1}^{n}((X_{i}-\mu)^{\top}u)^{2}\leq C_{\beta}(1+\max\{\sqrt{p/n},p^{2/\beta}/n\}),$
and hence with the same probability, we have
$\sup_{u\in\mathbb{S}^{p-1}}((\bar{X}-\mu)^{\top}u)\frac{1}{n}\sum_{i=1}^{n}((X_{i}-\mu)^{\top}u)^{2}$
$\leq$ $C_{\beta}\max\{\sqrt{p/n},(p/n)^{1/\beta}\}(1+\max\{\sqrt{p/n},p^{2/\beta}/n\})$
$\leq$ $C_{\beta}\max\{\sqrt{p/n},p^{3/\beta}/n\}$
and
$\sup_{u\in\mathbb{S}^{p-1}}((\bar{X}-\mu)^{\top}u)^{3}\leq C_{\beta}\max\{(p/n)^{3/2},(p/n)^{3/\beta}\}\leq C_{\beta}(p/n)^{3/\beta}\leq C_{\beta}p^{3/\beta}/n$
Thus, combining all, the desired bound follows. ∎
B.4 Proof of Theorem 5
Proof.
First, note that by Theorem 4, we have when $p^{2d/\beta-1}<n$, it follows
$\left\|\hat{\mathbfcal{M}}-{\mathbfcal{M}}\right\|\leq C_{\beta,d}\sqrt{\frac{p}{n}}.$
Furthermore, when $n<p$, we have $1<\max\left\{\frac{p^{3/\beta}}{n},\sqrt{\frac{p}{n}}\right\}$. Thus, it follows
$\left\|\hat{\mathbfcal{M}}-{\mathbfcal{M}}\right\|=\|\mathbb{E}(X_{i}^{\otimes d})\|\leq C_{\beta,d}.$
It remains to prove for $p<n<p^{2d/\beta-1}$. In this case, we have
$\left|\langle\hat{\mathbfcal{M}}-{\mathbfcal{M}},\bigotimes_{j\in[d]}u_{j}\rangle\right|$
$\leq$ $\left|\langle\mathbb{E}\hat{\mathbfcal{M}}-\mathbb{E}X_{i}^{\otimes d},\bigotimes_{j\in[d]}u_{j}\rangle\right|+\left|\langle\hat{\mathbfcal{M}}-\mathbb{E}\hat{\mathbfcal{M}},\bigotimes_{j\in[d]}u_{j}\rangle\right|.$
For the bias term, we have
$\left|\langle\mathbb{E}\hat{\mathbfcal{M}}-\mathbb{E}X_{i}^{\otimes d},\bigotimes_{j\in[d]}u_{j}\rangle\right|$ $=$ $\left|\langle\mathbb{E}\otimes_{j\in[d]}X_{i}I_{\{|X_{i}^{\top}u_{j}|\leq R\}}-\mathbb{E}X_{i}^{\otimes d},\bigotimes_{j\in[d]}u_{j}\rangle\right|$
$=$ $\left|\langle\mathbb{E}X_{i}^{\otimes d}(1-I_{\{|X_{i}^{\top}u_{j}|\leq R,j\in[d]\}}),\bigotimes_{j\in[d]}u_{j}\rangle\right|$
$=$ $\left|\mathbb{E}(1-I_{\{|X_{i}^{\top}u_{j}|\leq R,j\in[d]\}})\prod_{j\in[3]}|X_{i}^{\top}u_{j}|\right|$
$\leq$ $\sqrt{\mathbb{E}\prod_{j\in[d]}|X_{i}^{\top}u_{j}|^{2}\mathbb{E}(1-I_{\{|X_{i}^{\top}u_{j}|\leq R,j\in[d]\}})}$
$\leq$ $C_{\beta,d}\sqrt{\sum_{j\in[d]}\mathbb{P}\left(|X_{i}^{\top}u_{j}|>R\right)}$
$\leq$ $C_{\beta,d}\exp(-c_{\beta}R^{\beta}).$
For the variance term, let $Y_{ij}=X_{i}I_{\{|X_{i}^{\top}u_{j}|\leq R\}}$ and $Z_{i}=\prod_{j\in[d]}Y_{ij}^{\top}u_{j}$. Then we have
$Z_{i}\leq R^{d},$
and
$\left|\langle\hat{\mathbfcal{M}}-\mathbb{E}\hat{\mathbfcal{M}},\bigotimes_{j\in[d]}u_{j}\rangle\right|$ $=$ $\left|\frac{1}{n}\sum_{i=1}^{n}Z_{i}-\mathbb{E}Z_{i}\right|.$
This implies
$\mathbb{P}\left(\left|\frac{1}{n}\sum_{i=1}^{n}Z_{i}-\mathbb{E}Z_{i}\right|>t\right)\leq C\exp\left(-cnt^{2}/R^{2d}\right).$
Let $t=R^{d}\sqrt{p/n}$, it follows
$\mathbb{P}\left(\left|\langle\hat{\mathbfcal{M}}-\mathbb{E}\hat{\mathbfcal{M}},\bigotimes_{j\in[d]}u_{j}\rangle\right|>R^{d}\sqrt{p/n}\right)\leq C\exp\left(-cp\right).$
The desired result for $p<n<p^{2d/\beta-1}$ follows by combining the bound for bias and the variance.
B.5 Proof of Theorem 6
The first statement, i.e., the high probability bound for $\widehat{\mathbfcal{M}}$, immediately follows from the discussion in Section 3.3. The statement about expectation follows as $\|\widehat{\mathbfcal{M}}\|$ and $\|\mathbb{E}X^{\otimes d}\|$ are both upper bounded. Finally, for cumulant tensor estimation, by (1) and the fact that $F_{d}$ is a fixed tensor polynomial in the lower-order moments, with degree depending only on $d$, the results follow.
B.6 Proof of Lemma 1
We next calculate $\|L_{\leq D}\|^{2}:=\|L_{\leq D}\|^{2}_{L^{2}(H_{0})}$. We use the following notations: for vector $\alpha=(\alpha_{1},\ldots,\alpha_{m}),\alpha_{k}\in\mathbb{N}$, denote $|\alpha|=\sum_{k\in[m]}\alpha_{k}$, $\alpha!=\prod_{k\in[m]}\alpha_{k}!$, and multivariate Hermite polynomials $H_{\alpha}(x)=\prod_{k\in[m]}H_{\alpha_{i}}(x_{i})$. For vector $u\in\mathbb{R}^{m}$, let $u^{\alpha}=\prod_{k\in[m]}u_{k}^{\alpha_{k}}$. For a $np$-dimensional vector $y$, we use denote $y=(y^{(1)},\ldots,y^{(n)})^{\top}$ with $y^{(m)}\in\mathbb{R}^{p}$.
Denote $Y\in\mathbb{R}^{np}\sim\mathcal{P}$ under $H_{1}$ and $Y\in\mathbb{R}^{np}\sim\mathcal{Q}=N(0,I_{np})$ under $H_{0}$. Under $H_{1}$, we further denote $Y=(Y^{(1)},\ldots,Y^{(n)})^{\top}$ and $Y^{(i)}=S(\sqrt{a/p}W_{i}U+Z_{i})$. We let $V=(U,\ldots,U)\in\mathbb{R}^{np}$ and all entries of $U$ are i.i.d. Rademacher. Then we have (see, e.g., Lemma B.5 in [52] for detailed calculation)
$\|L_{\leq D}\|^{2}$ $=$ $\sum_{|\alpha|\leq D}\frac{(\mathbb{E}_{Y\sim\mathcal{P}}H_{\alpha}(Y))^{2}}{\alpha!}$ (28)
$=$ $\sum_{|\alpha|\leq D}\frac{1}{\alpha!}\left(\mathbb{E}_{U}\left(\prod_{i\in[n]}\mathbb{E}_{X|U}(H_{\alpha^{(i)}}(SX)|U)\right)\right)^{2}$
$=$ $\sum_{|\alpha|\leq D}\frac{\left(\prod_{i\in[n]}\left(\frac{a}{1+a}\right)^{|\alpha^{(i)}|/2}\mathbb{E}_{W_{1}}(H_{|\alpha^{(i)}|}(W_{1}))\right)^{2}}{\alpha!p^{|\alpha|}}\left(\mathbb{E}_{V}V^{\alpha}\right)^{2},$
where the last equality holds by Lemma B.11 in [52]. Note that we have
$\mathbb{E}H_{m}(W_{1})=0\quad\text{for $m\leq d-1$}.$ (29)
and
$\mathbb{E}_{V}V^{\alpha}=\mathbb{E}_{U}(\prod_{i\in[n]}\prod_{j\in[p]}u_{j}^{\alpha^{(i)}_{j}})=\prod_{j\in[p]}\mathbb{E}_{u_{j}}u_{j}^{\sum_{i\in[n]}\alpha^{(i)}_{j}}=I_{\{\sum_{i\in[n]}\alpha^{(i)}_{j}\text{ is even for all $j\in[p]$}\}},$ (30)
where $\alpha^{(i)}_{j}$ is the $j$th entry of vector $\alpha^{(i)}$ and $u_{i}$ is the $i$th entry of $U$.
Denote
$\mathcal{A}_{m}=\left\{\alpha\in\mathbb{N}^{np}:|\alpha|=m,|\alpha^{(j)}|=0\text{ or }|\alpha^{(j)}|\geq d,\sum_{i\in[n]}\alpha^{(i)}_{j}\text{ is even for all $j\in[p]$}\right\}.$
and
[…]
Thus, we have
$|\mathcal{A}_{m}|\leq|\mathcal{B}_{m}|=\binom{n}{\lfloor m/d\rfloor}\binom{p}{\lfloor m/2\rfloor}\binom{\lfloor m/2\rfloor\lfloor m/d\rfloor+m-1}{m}$.
Note that $m\leq D\ll\min{p,n}$. Thus, we have
$\binom{n}{k}\leq n^{k}\quad\text{ and }\quad\binom{\lfloor m/2\rfloor\lfloor m/d\rfloor+m-1}{m}\leq(m^{2}/8)^{m},$
which implies
$|\mathcal{A}_{m}|\leq n^{m/d}p^{m/2}(m^{2}/8)^{m}.$ (31)
Moreover, by Lemma 3, we have
$\sup_{k\leq m}\left(\mathbb{E}_{W_{1}}(H_{k}(W_{1}))\right)^{2m/k}\leq\sup_{k\leq m}(k!C_{d}^{k})^{2m/k}\leq C_{d}^{2m}m^{2m}.$ (32)
Finally, (28), (29), (30), (31), and (32) imply
$\|L_{\leq D}\|^{2}$ $=$ $\sum_{0\leq m\leq D}\sum_{\alpha\in\mathcal{A}_{m}}\frac{\left(\prod_{i\in[n]}\left(\frac{a}{1+a}\right)^{|\alpha^{(i)}|/2}\mathbb{E}_{W_{1}}(H_{|\alpha^{(i)}|}(W_{1}))\right)^{2}}{\alpha!p^{|\alpha|}}$
$\leq$ $\sum_{0\leq m\leq D}\sum_{\alpha\in\mathcal{A}_{m}}\frac{\prod_{i\in[n]}\left(\frac{a}{1+a}\right)^{|\alpha^{(i)}|}\left(\mathbb{E}_{W_{1}}(H_{|\alpha^{(i)}|}(W_{1}))\right)^{2}}{p^{m}}$
$=$ $\sum_{0\leq m\leq D}\sum_{\alpha\in\mathcal{A}_{m}}\left(\frac{a}{1+a}\right)^{m}\frac{\prod_{i\in[n]}\left(\mathbb{E}_{W_{1}}(H_{|\alpha^{(i)}|}(W_{1}))\right)^{2}}{p^{m}}$
$=$ $\sum_{0\leq m\leq D}\sum_{\alpha\in\mathcal{A}_{m}}\left(\frac{a}{1+a}\right)^{m}\frac{\prod_{i\in[n]}\left(\mathbb{E}_{W_{1}}(H_{|\alpha^{(i)}|}(W_{1}))\right)^{2|\alpha^{(i)}|/|\alpha^{(i)}|}}{p^{m}}$
$\leq$ $\sum_{0\leq m\leq D}\sum_{\alpha\in\mathcal{A}_{m}}\left(\frac{a}{1+a}\right)^{m}\frac{\prod_{i\in[n]}\left(\sup_{k\leq m}\left(\mathbb{E}_{W_{1}}(H_{k}(W_{1}))\right)^{2/k}\right)^{|\alpha^{(i)}|}}{p^{m}}$
$=$ $\sum_{0\leq m\leq D}\sum_{\alpha\in\mathcal{A}_{m}}\left(\frac{a}{1+a}\right)^{m}\frac{\sup_{k\leq m}\left(\mathbb{E}_{W_{1}}(H_{k}(W_{1}))\right)^{2m/k}}{p^{m}}$
$=$ $\sum_{0\leq m\leq D}\left(\frac{a}{1+a}\right)^{m}\frac{\sup_{k\leq m}\left(\mathbb{E}_{W_{1}}(H_{k}(W_{1}))\right)^{2m/k}}{p^{m}}|\mathcal{A}_{m}|$
$\leq$ $\sum_{0\leq m\leq D}\left(\frac{a}{1+a}\right)^{m}\frac{\sup_{k\leq m}\left(\mathbb{E}_{W_{1}}(H_{k}(W_{1}))\right)^{2m/k}}{p^{m}}n^{m/d}p^{m/2}(m^{2}/8)^{m}$
$\leq$ $\sum_{0\leq m\leq D}\left(\frac{a}{1+a}\right)^{m}\frac{C_{d}^{2m}m^{2m}}{p^{m}}n^{m/d}p^{m/2}(m^{2}/8)^{m}$
$=$ $\sum_{0\leq m\leq D}\left(\frac{a}{1+a}\frac{C_{d}m^{4}n^{1/d}}{p^{1/2}}\right)^{m}.$
B.7 Proof of Theorem 8
Proof.
The computational complexity is immediate from the definition of $V_{n}$: one first computes the Gram matrix
$G_{ij}=X_{i}^{\top}X_{j},$
and then averages the off-diagonal entries $G_{ij}^{d}$. This requires $O(n^{2}p)$ arithmetic operations.
We now analyze the mean and variance of $V_{n}$.
Step 1: Mean.
By independence of $X_{1}$ and $X_{2}$,
$\mathbb{E}\big[(X_{1}^{\top}X_{2})^{d}\big]$ $=\mathbb{E}\Big[\Big(\sum_{j=1}^{p}(X_{1})_{j}(X_{2})_{j}\Big)^{d}\Big]=\sum_{i_{1},\ldots,i_{d}\in[p]}\mathbb{E}\big[(X_{1})_{i_{1}}\cdots(X_{1})_{i_{d}}\big]^{2}=\|{\mathbfcal{M}}_{d}\|_{F}^{2}.$ (33)
Therefore,
$\mathbb{E}V_{n}=\|{\mathbfcal{M}}_{d}\|_{F}^{2}.$
In particular, under $H_{0}$,
$\mathbb{E}V_{n}=0,$
while under $H_{1}$,
$\mathbb{E}V_{n}=\|{\mathbfcal{M}}_{d}\|_{F}^{2}\geq\|{\mathbfcal{M}}_{d}\|^{2}\geq\kappa^{2}.$
Step 2: Hoeffding-type decomposition.
Set
$\theta:=\mathbb{E}[(X_{1}^{\top}X_{2})^{d}]=\|{\mathbfcal{M}}_{d}\|_{F}^{2}.$
For $x\in\mathbb{R}^{p}$, define
$g(x):=\mathbb{E}[(x^{\top}X_{2})^{d}]-\theta=\langle{\mathbfcal{M}}_{d},x^{\otimes d}\rangle-\|{\mathbfcal{M}}_{d}\|_{F}^{2}.$
Also define, for $x,y\in\mathbb{R}^{p}$,
$R(x,y):=(x^{\top}y)^{d}-\theta-g(x)-g(y).$
Then
$V_{n}-\theta=\frac{2}{n}\sum_{i=1}^{n}g(X_{i})+\frac{2}{n(n-1)}\sum_{1\leq i<j\leq n}R(X_{i},X_{j}).$
Moreover,
$\mathbb{E}[g(X_{1})]=0,$
and
$\mathbb{E}[R(X_{1},X_{2})\mid X_{1}]$ $=\mathbb{E}[(X_{1}^{\top}X_{2})^{d}\mid X_{1}]-\theta-g(X_{1})-\mathbb{E}[g(X_{2})]$
$=(\theta+g(X_{1}))-\theta-g(X_{1})-0$
$=0.$
Thus,
$\mathbb{E}[g(X_{1})]=0,\qquad\mathbb{E}[R(X_{1},X_{2})\mid X_{1}]=\mathbb{E}[R(X_{1},X_{2})\mid X_{2}]=0.$
Hence the two terms are orthogonal, and
$\text{Var}(V_{n})=\frac{4}{n}\text{Var}(g(X_{1}))+\frac{2}{n(n-1)}\text{Var}(R(X_{1},X_{2})).$ (34)
We now bound the two variance terms.
Step 3: Bound on the second-order remainder.
We first show that
$\mathbb{E}[(X_{1}^{\top}X_{2})^{2d}]\leq C_{d}K^{4d}p^{d}.$
Conditioning on $X_{1}$, the scalar $X_{1}^{\top}X_{2}$ is sub-Gaussian with
$\|X_{1}^{\top}X_{2}\|_{\Phi_{2}\,|\,X_{1}}\leq K\|X_{1}\|,$
hence
$\mathbb{E}[(X_{1}^{\top}X_{2})^{2d}\mid X_{1}]\leq C_{d}K^{2d}\|X_{1}\|^{2d}.$
Taking expectation gives
$\mathbb{E}[(X_{1}^{\top}X_{2})^{2d}]\leq C_{d}K^{2d}\mathbb{E}\|X\|^{2d}.$
To bound $\mathbb{E}\|X\|^{2d}$, let $g\sim N(0,I_{p})$ be independent of $X$. Since
$\mathbb{E}_{g}[(g^{\top}x)^{2d}]=c_{d}\|x\|^{2d},\qquad\forall x\in\mathbb{R}^{p},$
we have
$\mathbb{E}\|X\|^{2d}=c_{d}^{-1}\mathbb{E}[(g^{\top}X)^{2d}].$
Conditioning on $g$, the scalar $g^{\top}X$ is sub-Gaussian with
$\|g^{\top}X\|_{\Phi_{2}\,|\,g}\leq K\|g\|,$
so
$\mathbb{E}[(g^{\top}X)^{2d}\mid g]\leq C_{d}K^{2d}\|g\|^{2d}.$
Therefore
$\mathbb{E}\|X\|^{2d}\leq C_{d}K^{2d}\mathbb{E}\|g\|^{2d}\leq C_{d}K^{2d}p^{d},$
which implies
$\mathbb{E}[(X_{1}^{\top}X_{2})^{2d}]\leq C_{d}K^{4d}p^{d}.$ (35)
Next, since
$R(X_{1},X_{2})=(X_{1}^{\top}X_{2})^{d}-\theta-g(X_{1})-g(X_{2}),$
we have
$R(X_{1},X_{2})^{2}\leq 4\Big((X_{1}^{\top}X_{2})^{2d}+\theta^{2}+g(X_{1})^{2}+g(X_{2})^{2}\Big).$
Taking expectations and using Jensen’s inequality,
$\mathbb{E}[g(X_{1})^{2}]\leq\mathbb{E}[(X_{1}^{\top}X_{2})^{2d}],\qquad\theta^{2}\leq\mathbb{E}[(X_{1}^{\top}X_{2})^{2d}],$
we obtain
$\mathbb{E}[R(X_{1},X_{2})^{2}]\leq 16\,\mathbb{E}[(X_{1}^{\top}X_{2})^{2d}]\leq C_{d}K^{4d}p^{d}.$
Hence
$\text{Var}(R(X_{1},X_{2}))\leq\mathbb{E}[R(X_{1},X_{2})^{2}]\leq C_{d}K^{4d}p^{d}.$ (36)
Step 4: Type-I error under $H_{0}$.
Under $H_{0}$, we have ${\mathbfcal{M}}_{d}=0$, so by definition of $g$,
$g(X_{1})=\langle{\mathbfcal{M}}_{d},X_{1}^{\otimes d}\rangle-\|{\mathbfcal{M}}_{d}\|_{F}^{2}=0.$
Thus the first variance term in (34) vanishes, and by (36),
$\text{Var}(V_{n})=\frac{2}{n(n-1)}\text{Var}(R(X_{1},X_{2}))\leq C_{d}K^{4d}\frac{p^{d}}{n(n-1)}.$
Since $\mathbb{E}V_{n}=0$, Chebyshev’s inequality yields
$\mathbb{P}_{H_{0}}(V_{n}>\tau_{n})\leq\frac{\text{Var}(V_{n})}{\tau_{n}^{2}}\leq\frac{C_{d}K^{4d}p^{d}/n^{2}}{A^{2}p^{d}(\log n)^{2}/n^{2}}=\frac{C_{d}K^{4d}}{A^{2}(\log n)^{2}}\to 0.$
Therefore
$\mathbb{P}_{H_{0}}(\phi_{n}=1)\to 0.$
Step 5: Type-II error under $H_{1}$.
Under $H_{1}$, the assumption (5) gives
$\text{Var}(g(X_{1}))=\text{Var}\!\big(\langle{\mathbfcal{M}}_{d},X_{1}^{\otimes d}\rangle-\|{\mathbfcal{M}}_{d}\|_{F}^{2}\big)=\text{Var}\!\big(\langle{\mathbfcal{M}}_{d},X_{1}^{\otimes d}\rangle\big)\leq C_{1}p^{d/2}.$
Combining this with (34) and (36), we obtain
$\text{Var}(V_{n})\leq C\frac{p^{d/2}}{n}+C_{d}K^{4d}\frac{p^{d}}{n(n-1)}.$
Hence, whenever
$\frac{p^{d/2}\log n}{n}\to 0,$
we have in particular
$\text{Var}(V_{n})\to 0.$
On the other hand, under $H_{1}$,
$\mathbb{E}V_{n}=\|{\mathbfcal{M}}_{d}\|_{F}^{2}\geq\kappa^{2}.$
Since $\tau_{n}\to 0$, for all sufficiently large $n$ we have
$\tau_{n}\leq\frac{\kappa^{2}}{2}.$
Therefore
$\mathbb{P}_{H_{1}}(\phi_{n}=0)$ $=\mathbb{P}_{H_{1}}(V_{n}\leq\tau_{n})$
$\leq\mathbb{P}_{H_{1}}\Big(V_{n}-\mathbb{E}V_{n}\leq-(\mathbb{E}V_{n}-\tau_{n})\Big)$
$\leq\mathbb{P}_{H_{1}}\Big(|V_{n}-\mathbb{E}V_{n}|\geq\frac{\kappa^{2}}{2}\Big)$
$\leq\frac{4\,\text{Var}(V_{n})}{\kappa^{4}}\to 0$
by Chebyshev’s inequality. This proves
$\mathbb{P}_{H_{1}}(\phi_{n}=0)\to 0.$
B.8 Proof of Theorem 9
We first compute the mean of $V_{n}^{(\mathrm{cumu})}$. By independence of $Y_{1}$ and $Y_{2}$,
$\mathbb{E}V_{n}^{(\mathrm{cumu})}$ $=\mathbb{E}\big[\langle\Psi_{d}(Y_{1}),\Psi_{d}(Y_{2})\rangle\big]$
$=\Big\langle\mathbb{E}[\Psi_{d}(Y_{1})],\mathbb{E}[\Psi_{d}(Y_{2})]\Big\rangle$
$=\|{\mathbfcal{K}}_{d}\|_{F}^{2}.$ (37)
Therefore, under $H_{0}$,
$\mathbb{E}V_{n}^{(\mathrm{cumu})}=0,$
while under $H_{1}$,
$\mathbb{E}V_{n}^{(\mathrm{cumu})}=\|{\mathbfcal{K}}_{d}\|_{F}^{2}\geq\|{\mathbfcal{K}}_{d}\|^{2}\geq\kappa^{2}.$
We next perform the Hoeffding decomposition. Set
$h(y_{1},y_{2}):=\langle\Psi_{d}(y_{1}),\Psi_{d}(y_{2})\rangle,\qquad\theta:=\mathbb{E}[h(Y_{1},Y_{2})]=\|{\mathbfcal{K}}_{d}\|_{F}^{2}.$
For $y\in\mathbb{R}^{p}$, define
$g(y):=\mathbb{E}[h(y,Y_{2})]-\theta=\langle\Psi_{d}(y),{\mathbfcal{K}}_{d}\rangle-\|{\mathbfcal{K}}_{d}\|_{F}^{2},$
and for $y_{1},y_{2}\in\mathbb{R}^{p}$, define
$R(y_{1},y_{2}):=h(y_{1},y_{2})-\theta-g(y_{1})-g(y_{2}).$
Then
$V_{n}^{(\mathrm{cumu})}-\theta=\frac{2}{n}\sum_{i=1}^{n}g(Y_{i})+\frac{2}{n(n-1)}\sum_{1\leq i<j\leq n}R(Y_{i},Y_{j}).$
Also,
$\mathbb{E}[g(Y_{1})]=0.$
Moreover,
$\mathbb{E}[h(Y_{1},Y_{2})\mid Y_{1}]=\langle\Psi_{d}(Y_{1}),\mathbb{E}[\Psi_{d}(Y_{2})]\rangle=\langle\Psi_{d}(Y_{1}),{\mathbfcal{K}}_{d}\rangle=\theta+g(Y_{1}),$
hence
$\mathbb{E}[R(Y_{1},Y_{2})\mid Y_{1}]=(\theta+g(Y_{1}))-\theta-g(Y_{1})-\mathbb{E}[g(Y_{2})]=0.$
Similarly,
$\mathbb{E}[R(Y_{1},Y_{2})\mid Y_{2}]=0.$
Therefore, the linear part and the degenerate part are orthogonal, and the standard variance formula for order-two U-statistics yields
$\text{Var}\big(V_{n}^{(\mathrm{cumu})}\big)=\frac{4}{n}\text{Var}\big(g(Y_{1})\big)+\frac{2}{n(n-1)}\text{Var}\big(R(Y_{1},Y_{2})\big).$ (38)
We now bound the degenerate term. Since
$R(Y_{1},Y_{2})=h(Y_{1},Y_{2})-\theta-g(Y_{1})-g(Y_{2}),$
we have
$R(Y_{1},Y_{2})^{2}\leq 4\Big(h(Y_{1},Y_{2})^{2}+\theta^{2}+g(Y_{1})^{2}+g(Y_{2})^{2}\Big).$
By Jensen’s inequality,
$\mathbb{E}[g(Y_{1})^{2}]=\mathbb{E}\Big[\big(\mathbb{E}[h(Y_{1},Y_{2})\mid Y_{1}]-\theta\big)^{2}\Big]\leq\mathbb{E}[h(Y_{1},Y_{2})^{2}],$
and similarly
$\theta^{2}=\big(\mathbb{E}[h(Y_{1},Y_{2})]\big)^{2}\leq\mathbb{E}[h(Y_{1},Y_{2})^{2}].$
Consequently,
$\mathbb{E}[R(Y_{1},Y_{2})^{2}]\leq 16\,\mathbb{E}[h(Y_{1},Y_{2})^{2}]\leq 16C_{0}p^{d}$
by (9). Hence
$\text{Var}\big(R(Y_{1},Y_{2})\big)\leq\mathbb{E}[R(Y_{1},Y_{2})^{2}]\leq 16C_{0}p^{d}.$ (39)
We first analyze the type-I error. Under $H_{0}$, we have ${\mathbfcal{K}}_{d}=0$, so by the definition of $g$,
$g(Y_{1})=\langle\Psi_{d}(Y_{1}),{\mathbfcal{K}}_{d}\rangle-\|{\mathbfcal{K}}_{d}\|_{F}^{2}=0.$
Thus the first term in (38) vanishes, and (39) gives
$\text{Var}\big(V_{n}^{(\mathrm{cumu})}\big)=\frac{2}{n(n-1)}\text{Var}\big(R(Y_{1},Y_{2})\big)\leq C\frac{p^{d}}{n(n-1)}.$
Since $\mathbb{E}V_{n}^{(\mathrm{cumu})}=0$, Chebyshev’s inequality implies
$\mathbb{P}_{H_{0}}\big(V_{n}^{(\mathrm{cumu})}>\tau_{n}\big)\leq\frac{\text{Var}(V_{n}^{(\mathrm{cumu})})}{\tau_{n}^{2}}\leq\frac{Cp^{d}/n^{2}}{A^{2}p^{d}(\log n)^{2}/n^{2}}=\frac{C}{A^{2}(\log n)^{2}}\to 0.$
Therefore,
$\mathbb{P}_{H_{0}}\big(\phi_{n}^{(\mathrm{cumu})}=1\big)\to 0.$
We next analyze the type-II error. Under $H_{1}$, assumption (10) implies
$\text{Var}\big(g(Y_{1})\big)=\text{Var}\big(\langle\Psi_{d}(Y_{1}),{\mathbfcal{K}}_{d}\rangle-\|{\mathbfcal{K}}_{d}\|_{F}^{2}\big)=\text{Var}\big(\langle\Psi_{d}(Y_{1}),{\mathbfcal{K}}_{d}\rangle\big)\leq C_{1}p^{d/2}.$
Combining this with (38) and (39), we obtain
$\text{Var}\big(V_{n}^{(\mathrm{cumu})}\big)\leq C\frac{p^{d/2}}{n}+C\frac{p^{d}}{n(n-1)}.$
Hence, whenever
$\frac{p^{d/2}\log n}{n}\to 0,$
we have in particular
$\text{Var}\big(V_{n}^{(\mathrm{cumu})}\big)\to 0.$
On the other hand, under $H_{1}$,
$\mathbb{E}V_{n}^{(\mathrm{cumu})}=\|{\mathbfcal{K}}_{d}\|_{F}^{2}\geq\kappa^{2}.$
Since $\tau_{n}\to 0$, for all sufficiently large $n$,
$\tau_{n}\leq\frac{\kappa^{2}}{2}.$
Therefore,
$\mathbb{P}_{H_{1}}\big(\phi_{n}^{(\mathrm{cumu})}=0\big)$ $=\mathbb{P}_{H_{1}}\big(V_{n}^{(\mathrm{cumu})}\leq\tau_{n}\big)$
$\leq\mathbb{P}_{H_{1}}\Big(V_{n}^{(\mathrm{cumu})}-\mathbb{E}V_{n}^{(\mathrm{cumu})}\leq-\big(\mathbb{E}V_{n}^{(\mathrm{cumu})}-\tau_{n}\big)\Big)$
$\leq\mathbb{P}_{H_{1}}\Big(\big|V_{n}^{(\mathrm{cumu})}-\mathbb{E}V_{n}^{(\mathrm{cumu})}\big|\geq\frac{\kappa^{2}}{2}\Big)$
$\leq\frac{4\,\text{Var}(V_{n}^{(\mathrm{cumu})})}{\kappa^{4}}\to 0$
by Chebyshev’s inequality. This proves
$\mathbb{P}_{H_{1}}\big(\phi_{n}^{(\mathrm{cumu})}=0\big)\to 0.$
Finally, because the lower-order cumulants are known, the map $\Psi_{d}$ is known. For fixed $d$, if each evaluation of $\Psi_{d}(y)$ requires $O(p^{d})$ arithmetic operations, then each kernel evaluation
$\langle\Psi_{d}(Y_{i}),\Psi_{d}(Y_{j})\rangle$
can be computed in $O(p^{d})$ operations, and therefore the full statistic $V_{n}^{(\mathrm{cumu})}$ can be computed in $O(n^{2}p^{d})$ operations.
Appendix C Lemmas
Lemma 3.
There exists a sub-Gaussian random variable $X\in\mathbb{R}$, such that its first d1d-1 cumulants are Gaussian cumulants while the $d$th cumulant is some non-zero constant $c_{d}$. Moreover, we have $\mathbb{E}H_{m}(X)\leq m!C_{d}^{m}$ for all $m\in\mathbb{N}^{+}$, where $H_{m}(t)$ be the order-$m$ probabilists’ Hermite polynomial.
Proof.
Let $\psi_{m}(t)=\operatorname{sign}(H_{m}(t))\min\{|H_{m}(t)|,M\}$. Define $X_{\beta,\theta}$ by its density
$f(x)=\frac{\exp\left(\sum_{i\in[d-1]}\beta_{i}\psi_{i}(x)+\theta\psi_{d}(x)\right)\phi(x)}{C_{\beta,\theta}},$
where $\beta\in\mathbb{R}^{d-1}$, $\theta\in\mathbb{R}$, and $\phi(x)$ is the density of standard normal distribution.
When $\theta=0,\beta=0$, $X_{\beta,\theta}\sim N(0,1)$. Denote
$F(\beta,\theta)=(\mathbb{E}H_{1}(X_{\beta,\theta}),\ldots,\mathbb{E}H_{d-1}(X_{\beta,\theta}))^{\top}.$
By implicit function theorem and the fact that the Jacobian is diagonal with $i$th entry close to $i!$ for large $M$, it can be prove that there exist $M_{d}$, $\delta_{d}$ and map $\theta\mapsto\beta(\theta)$ with $\beta(0)=0$ such that $F(\beta(\theta),\theta)=0$ for all $|\theta|\leq\delta_{d}$ and $M\geq M_{d}$.
Thus, for all $\theta$ that is sufficiently small, we can always pick $\beta$ such that $\mathbb{E}H_{i}(X_{\beta,\theta})=0$ for all $i\in[d-1]$, which further implies that the cumulants match the Gaussian cumulants up to order d1d-1.
The remaining is to pick $\theta$ such that the $d$th cumulant of $X_{\theta}=:X_{\beta(\theta),\theta}$ is not zero. Note that we have the $d$th cumulant $k_{d}(X_{\theta})=\mathbb{E}H_{d}(X_{\theta})$ for $d\geq 3$.
Let $G(\beta,\theta)=\mathbb{E}H_{d}(X_{\beta,\theta})$ and $g(\theta)=\mathbb{E}H_{d}(X_{\theta})$.
Note that $g(0)=0$ and differentiating at 0 with the chain rule gives
$g^{\prime}(0)=\partial_{\theta}G(0,0)+\sum_{j=1}^{d-1}\partial_{\beta_{j}}G(0,0)\,[\beta^{\prime}(0)]_{j}.$
The partial derivatives can be written as covariances:
$\partial_{\theta}G(0,0)$ $=\mathrm{Cov}\big(H_{d}(Z),\psi_{d}(Z)\big),$
$\partial_{\beta_{j}}G(0,0)$ $=\mathrm{Cov}\big(H_{d}(Z),\psi_{j}(Z)\big).$
Moreover, $\beta^{\prime}(0)$ is determined by differentiating the identity $F(\beta(\theta),\theta)=0$ at $\theta=0$:
$J_{\beta}\,\beta^{\prime}(0)+F_{\theta}(0,0)=0,$
where $F_{\theta}(0,0)$ has entries $\mathrm{Cov}(H_{i}(Z),\psi_{d}(Z))$. Hence
$\beta^{\prime}(0)=-\big(J_{\beta}\big)^{-1}F_{\theta}(0,0).$
Using again that $\psi_{m}\to H_{m}$ as $M\to\infty$, we obtain
$J_{\beta}$ $\to\mathrm{diag}(1!,2!,\dots,(d-1)!),$
$F_{\theta}(0,0)$ $\to 0,$
$\partial_{\theta}G(0,0)$ $\to\mathbb{E}H_{d}(Z)^{2}=d!,$
$\partial_{\beta_{j}}G(0,0)$ $\to 0,\quad j=1,\dots,d-1.$
Consequently,
$g^{\prime}(0)=\partial_{\theta}G(0,0)+\sum_{j=1}^{d-1}\partial_{\beta_{j}}G(0,0)\,[\beta^{\prime}(0)]_{j}\rightarrow d!.$
In particular, for all sufficiently large $M$ depending on $d$, we have $g^{\prime}(0)\neq 0$. Thus, we can pick $\theta_{0}$ so that $g_{M}(\theta_{0})\neq 0$.
For $X:=X_{\theta_{0}}=X_{\beta(\theta_{0}),\theta_{0}}$, we have for $m\leq d-1$, $\kappa_{m}(X)$ equals the Gaussian cumulant of order $m$; and $\kappa_{d}(X)=g(\theta_{0})\neq 0$.
For this fixed $M$ and $(\beta,\theta_{0})$, the tilt
$T(x):=\sum_{i=1}^{d-1}\beta_{i}\psi_{i}(x)+\theta_{0}\psi_{d}(x)$
is bounded, say $|T(x)|\leq B_{d}$ for all $x$. Hence there exist constants $0<c_{d}\leq C_{d}<\infty$ (depending only on $d$) such that
$c_{d}\phi(x)\leq f_{X}(x)\leq C_{d}\phi(x)\quad\text{for all }x\in\mathbb{R}.$
This implies in particular that $X$ is sub-Gaussian with a parameter depending only on $d$.
The probabilists’ Hermite polynomials satisfy the explicit formula
$H_{m}(x)=m!\sum_{k=0}^{\lfloor m/2\rfloor}\frac{(-1)^{k}}{k!(m-2k)!2^{k}}\,x^{m-2k},$
from which one obtains a standard growth bound of the form
$|H_{m}(x)|\leq m!(C_{0}(1+|x|))^{m},\qquad\forall m\geq 1,$
for some absolute constant $C_{0}>0$. Using Holder’s inequality together with the sub-Gaussian tail bound for $X$, there exists a constant $C_{d}\geq 1$ such that
$\mathbb{E}|H_{m}(X)|\leq m!C_{d}^{m},\qquad\forall m\in\mathbb{N}.$
This completes the proof. ∎
Lemma 4.
For unit vectors $u,v\in\mathbb{R}^{p}$ with $\langle u,v\rangle>0$, we have $\|u^{\otimes d}-v^{\otimes d}\|\geq\|u-v\|/2$.
Proof.
Let
$\alpha:=\langle u,v\rangle,\quad\delta:=\|u-v\|=\sqrt{2-2\alpha},\quad w:=(u-v)/\delta.$
So we have
$u^{\top}w=-v^{\top}w=\delta/2.$
Choose the $d$ unit vectors
$x_{1}=\cdots=x_{d-1}:=u,\quad x_{d}:=w.$
Then we have
$\|u^{\otimes d}-v^{\otimes d}\|\geq\langle u^{\otimes d}-v^{\otimes d},\bigotimes_{k=1}^{d}x_{k}\rangle=\delta/2(1+\alpha^{d-1})\geq\delta/2.$
Lemma 5 (Theorem 2 in [20]).
For any symmetric tensor, there exists a best rank-one approximation that is symmetric.
Lemma 6 (Corollary 1.4 in [21]).
Let $X_{1},\ldots,X_{n}$ be a set of independent, centered random variables with $\left\|X_{i}\right\|_{\Psi_{\alpha}}\leq M$ for some $\alpha\in(0,1]$. Let $a\in\mathbb{R}^{n}$. For any $t\geq 0$ it holds
$\mathbb{P}\left(\left|\sum_{i=1}^{n}a_{i}X_{i}\right|\geq t\right)\leq 2\exp\left(-\frac{1}{C_{\alpha}}\min\left(\frac{t^{2}}{M^{2}\|a\|^{2}},\frac{t^{\alpha}}{M^{\alpha}\max_{i}\left|a_{i}\right|^{\alpha}}\right)\right).$
Lemma 7 (Generalized Fano’s Lemma (e.g., Lemma 3 in [59])).
Let $r\geq 2$ be an integer and let $\mathcal{M}_{r}\subset\mathcal{P}$ contain $r$ probability measures indexed by $j=1,2,\ldots,r$ such that for all $j\neq j^{\prime}$
$d\left(\theta\left(P_{j}\right),\theta\left(P_{j^{\prime}}\right)\right)\geq\alpha_{r},$
and
$K\left(P_{j},P_{j^{\prime}}\right)=\int\log\left(P_{j}/P_{j^{\prime}}\right)dP_{j}\leq\beta_{r}.$
Then
$\max_{j}E_{P_{j}}d\left(\hat{\theta},\theta\left(P_{j}\right)\right)\geq\frac{\alpha_{r}}{2}\left(1-\frac{n\beta_{r}+\log 2}{\log r}\right),$
for any $\hat{\theta}$ as the estimator based on $n$ i.i.d. realizations.
Lemma 8 (Upper bound of KL divergence of Gaussian mixture model (e.g., Lemma 1 in [14])).
For two Gaussian mixtures with the same number of components: $f(x)=\sum_{a\in[d]}\pi_{a}\mathcal{N}\left(x;\mu_{a};\Sigma_{a}\right)$ and $g(x)=\sum_{b\in[d]}\omega_{b}\mathcal{N}\left(x;\mu_{b};\Sigma_{b}\right)$, we have
$D(f\|g)\leq\sum_{a\in[d]}\pi_{a}\left(D\left(f_{a}\|g_{a}\right)+\log\pi_{a}-\log\omega_{a}\right).$
Lemma 9 (Upper bound of KL divergence of Gaussian models).
For two Gaussian distributions $f(x)=\mathcal{N}\left(\mu_{0};\Sigma_{0}\right)$ and $g(x)=\mathcal{N}\left(\mu_{0};\Sigma_{1}\right)$, if $\Sigma_{i}=I-A_{i}\in\mathbb{R}^{k\times k}$ with some symmetric matrix $A_{i}$ such that $\text{tr}(A_{i})=0$ and $\|A_{i}\|_{F}\leq\delta<c$ for some absolute constant $c$, then we have
$D(f\|g)\leq C\delta^{2}.$
Proof.
By definition, we have
$D(f\|g)$ $=$ $\frac{1}{2}\left[\operatorname{tr}\left(\Sigma_{1}^{-1}\Sigma_{0}\right)-k+\log\frac{\operatorname{det}\Sigma_{1}}{\operatorname{det}\Sigma_{0}}\right].$
First, note that
$\Sigma_{1}^{-1}=(I-A_{1})^{-1}=I+A_{1}+A_{1}^{2}+\ldots.$
Hence, we have
$\Sigma_{0}\Sigma_{1}^{-1}=I+A_{1}-A_{0}+(A_{1}^{2}-A_{0}A_{1})+\ldots.$
Note that $\text{tr}(A_{i})=0$, $\text{tr}(I)=k$, and $\text{tr}(A_{0}^{p}A_{1}^{q})\leq\|A_{0}\|_{F}^{p}\|A_{1}\|_{F}^{q}\leq(\delta)^{p+q}$. Thus, it follows
$\text{tr}(\Sigma_{0}\Sigma_{1}^{-1})-k\leq C\delta^{2}.$
Moreover, we have
$\det(\Sigma_{1})=\det(I-A_{1})$
$=$ $1-\text{tr}(A_{1})+\frac{1}{2}((\text{tr}(A_{1}))^{2}-\text{tr}(A_{1}^{2}))+\frac{1}{6}(\text{tr}(A_{1}^{3})-3\text{tr}(A_{1})\text{tr}(A_{1}^{2})+2(\text{tr}(A_{1}))^{3})+\ldots.$
By $\text{tr}(A_{1})=0$ and $\text{tr}(A_{1}^{q})\leq(\delta)^{q}$, we have
$\det(\Sigma_{1})\leq 1+C\delta^{2}.$
Similarly, we also have
$\det(\Sigma_{0})\geq 1-C\delta^{2}.$
Thus, it follows
$\log\left(\frac{\det(\Sigma_{1})}{\det(\Sigma_{0})}\right)\leq\log\left(1+\frac{2C\delta^{2}}{1-C\delta^{2}}\right)\leq\frac{2C\delta^{2}}{1-C\delta^{2}}\leq C\delta^{2}.$
Combining all the above, we have
$D(f\|g)\leq C\delta^{2}.$